### v1

# : Towards Semantic Steganography via Large Language Models

Abstract
Although steganography has made significant advancements in recent years, it still struggles to embed semantically rich, sentence-level information into carriers. However, in the era of AIGC, the capacity of steganography is more critical than ever. In this work, we present Sentence-to-Image Steganography, an instance of Semantic Steganography, a novel task that enables the hiding of arbitrary sentence-level messages within a cover image. Furthermore, we establish a benchmark named Invisible Text (IVT), comprising a diverse set of sentence-level texts as secret messages for evaluation. Finally, we present : Semantic Steganographic Language Model, which utilizes large language models (LLMs) to embed high-level textual informationβsuch as sentences or even paragraphsβinto images. Unlike traditional bit-level counterparts, enables the integration of semantically rich content through a newly designed pipeline in which the LLM is involved throughout the entire process. Both quantitative and qualitative experiments demonstrate that our method effectively unlocks new semantic steganographic capabilities for LLMs. The source code will be released soon.
1 Introduction
In an age long past, under the shadow of the Persian Empire, Histiaeus, ruler of Miletus, devised a plan that would etch his name in the annals of history. Under the watch of the imperial, he tattooed a secret message onto the scalp of a loyal slave, waiting for the hair to regrow and conceal the message beneath a veil of innocence. When the time was right, the slave crossed enemy lines, carrying the message not in his hands but under his skin. The message sparked a revolution, a call to arms that would ignite the Ionian Revolt. This tale, immortalized by Herodotus in the Histories, remains one of the earliest recorded feats of Steganography - the art of hiding information in plain sight.
Fast forward today, on the verge of entering the Artificial Intelligence Generated Content (AIGC) era, the principles of steganography have never been more relevant, with the prosperity of various generative models (e.g., Stable Diffusion [rombach2021highresolution], ChatGPT [openai_chatgpt], and Llama [touvron2023llama]), which can produce highly realistic text [bai2023qwentechnicalreport, qwen2025qwen25technicalreport], images [rombach2021highresolution, ramesh2022hierarchicaltextconditionalimagegeneration] and videos [liu2024sorareviewbackgroundtechnology], and even execute tasks with human-like proficiency [durante2024agentaisurveyinghorizons]. Their widespread application has blurred the line between fact and forgery, raising numerous security and legal concerns. The application of the generative model may foster the generation of biased content [ling2025bias, serouis2024exploring, kumar2025no], image manipulation [qu2024towards], and the creation of fake news [wu2024fake]. Moreover, the human-like content generated by those models is flooding the entire Internet, seriously hindering the determination of the credibility of evidence in forensics [yin2025digital]. Therefore, there is an urgent need for effective measures to prevent the large-scale misuse of generative models.
To address this problem, embedding forensics information into the generated content by steganography could be the most effective method. As the real-world application of steganography, embeds secret information that is typically difficult to detect and remove but can be easily retrieved [hartung2002multimedia]. With this technique, it is possible to determine whether the generative model creates the content quickly. Moreover, the extracted information can also be used for copyright identification and source tracing, thereby protecting intellectual property rights and helping locate the infringer. Nowadays, various strategies have been developed for text [pan2024markllm, zhang2024remark, liang2024watermarkingtechniqueslargelanguage], images [wen2023tree, fernandez2023stable], video [lin2024automatic, luo2023dvmark], and audio [zhang2024v2a] generated by generative models.
As a primary medium of visual communication, images are widely shared on the Internet, promoting the development of digital image steganography. Over years of research, numerous methods have been proposed[zhu2018hidden, wengrowski2019light, tancik2020stegastamp, wen2023tree, cox2007digital, wang2001image, patel2012steganography]. However, most existing approaches focus on embedding low-level binary or character-based information, neglecting the semantic relationships inherent in natural language. Consequently, these methods exhibit limited capacity and poor expressivity, preventing them from embedding sentence-level messages. Those methods underutilize the representational potential of modern high-resolution images and fail to meet the increasing capacity demands of watermarking in the era of generative AI.
For further analysis, we divide the current steganography [zhu2018hidden, wengrowski2019light, tancik2020stegastamp, patel2012steganography] pipeline into three processes: the encoding process, which transforms the secret information into a suitable representation; the insertion process, which embeds the message into the cover image to produce a stego image; and the decoding process, which extracts the secret message from the stego image. As illustrated in Figure 1, three significant challenges hinder current steganographic methods from embedding substantial semantic information into images: 1) In the encoding process, secret messages are typically converted into bitstreams based on ASCII or UTF formats. This method is simple but primitive, neglects the semantic information between words, has a low information density, and is unsuitable for encoding large information such as sentences; 2) In the insertion process, the encoder takes the secret message and cover image as input to generate a stego image. The interaction between two fundamentally different modalities (bitstreams and images) results in suboptimal insertion performance. 3) In the decoding process, due to the aforementioned limitations, the decoder can only recover bitstreams, restricting the hidden information to the word level. The three significant challenges mentioned above hinder the existing steganography to increase the secret message capacity to include semantic information, such as complete sentences.
To address these challenges, we introduce the broader concept of Semantic Steganography, which focuses on embedding semantically meaningful and structured contentβsuch as sentences or paragraphsβinto carriers. Building on this idea, we formalize the Sentence-to-Image Steganography task as an instance of semantic steganography. This task requires hiding sentence-level or even paragraph-level semantic information into a cover image. Notably, the secret message can consist of arbitrary textual content of any length. To accomplish this task, the model must possess two key capabilities: 1) encoding large and complex information into an appropriate representation and 2) effectively handling two fundamentally different modalities (text and image).
Recent advancements in the development of large language models (LLMs) have spurred efforts to integrate LLMs with vision backbones, leading to the rise of vision-language models (VLMs) [zhang2024vision, ye2024mplug, li2022mplug, liu2024improved, liu2023visual]. However, most of these models focus on vision understanding and text generation. This raises an important question: Can LLMs be used directly as a universal steganographic tool to embed sentence-level information into carriers such as images?
To this end, we propose the Semantic Steganographic Language Model (), an innovative steganographic framework. To the best of our knowledge, this is the first work to explore hiding sentence-level and even paragraph-level information in images. Specifically, we redesign the entire pipeline used in current watermarking strategies. As illustrated in FigureΒ 1, we first encode the secret sentence into tokens using a tokenizer. Second, during the insertion process, we use the LLM to generate secret embeddings without relying on image priors, thereby eliminating interference between the cover image and the secret message. Finally, we use the same LLM to decode the secret message from the stego image.
In our empirical experiments, can conceal up to 500 words of text into a single image with a resolution of , breaking the boundary between image and text. Additionally, we evaluate the steganographic performance of different LLMs within the framework and verify its versatility across various language models.
Additionally, we establish a benchmark for semantic steganography called Invisible Text (IVT), which we use to validate the effectiveness of the sentence-to-image framework in this paper. Overall, our contributions are as follows:
β’
We propose a new steganographic taskβSentence-to-Image Steganography, which requires embedding sentence-level information into a cover image.
β’
We propose the framework, incorporating new capabilities to embed sentence-level information into images. It demonstrates the potential of LLMs in steganography, achieving a significant increase in capacity compared to traditional methods.
β’
We establish a benchmark for the semantic steganography named Invisible Text (IVT), containing thousands of sentence-level samples from diverse sources. This benchmark is essential for evaluation and encourages the community to further explore the potential of semantic steganography.
2 Related Works
2.1 Image Steganography
Steganography is a technique for concealing secret information within a carrier while preserving its visual appearance and avoiding detection. Traditional image steganography methods, such as the Least Significant Bit (LSB) algorithm [wang2001image] and frequency-domain approaches [patel2012steganography, chen2006dwt] (e.g., DCT and DWT), embed secret data by modifying the pixel values or transform coefficients of the cover image. However, these methods have limited capacity and struggle to meet the requirements of modern steganography, whereas deep learning-based methods perform better. Recent advances in deep learning have produced more powerful steganographic techniques, such as HiDDeN [zhu2018hidden], LFM [wengrowski2019light], and StegaStamp [tancik2020stegastamp], which can embed bit-level information in images with robustness against various manipulations. Moreover, methods like HiNet [jing2021hinet], ISN [lu2021large], LanNet [lan2023robust], and StegFormer [ke2024stegformer] offer greater capacity, enabling the concealment of an image in carrier with the same resolution.
Despite these advancements, existing methods primarily focus on embedding secret information at a low level (bits or images). In this paper, we present a novel steganographic task, Sentence-to-Image Steganography, which aims to conceal semantically meaningful content (e.g., sentence) into images. To this end, we propose , successfully embedding sentence-level textual data into cover images, significantly expanding the scope of steganography.
2.2 Large Language Models in Computer Vision
The prosperity of the current large language model (LLM) has pushed the whole community of the Natural Language Processing (NLP) into new boundaries. The computer vision community has also noticed the tremendous success of the practice of LLM. Consequently, lots of the latest research focuses on exploring the integrity of the LLM with the vision backbone. BLIP [li2022blip] proposes a Multi-modal mixture of Encoder-Decoder (MED), while BLIP2 [li2023blip] introduces a frozen LLM to bootstrap vision-to-language learning, and proposes a Q-Former to bridge the modality gap. LLaVA [liu2023visual] connects the vision backbone and Llama [touvron2023llama] and tunes the LLM with multi-modal language-image instruction-following data, achieving the general-purpose visual and language understanding.
The methods above explore the LLM application in visual understanding, often seen as a high-level vision task. However, the LLM can also enhance the performance of the low-level vision backbone. LISA [Lai2023LISARS] uses a novel segmentation token and unlocks the segmentation capability of the LLM, which can be used for reasoning segmentation tasks. Xu et al. used localized image captioning data to pretrain PixelLLM to equip it with localization capability [xu2024pixel]. Zheng et al. froze LLM to solve a range of low-level vision tasks, demonstrating the LLMβs strong generalization capability [zheng2024lm4lv].
In this paper, we investigate the steganographic capabilities of LLMs, which has not been explored in the previous works. Experiments show that LLMs can serve as a steganographic tool to embed sentence-level information into a cover image.
3 Sentence-to-Image Steganography
The whole steganographic community has traditionally focused on low-level information hiding (e.g., at the bit or image level), overlooking the semantic structure of the hidden information. In contrast, we introduce a new direction of Semantic Steganography, which refers to hiding semantically meaningful and structured content, such as sentences or paragraphs, in the carrier. This paradigm redefines what kind of information can be hidden. Rather than merely embedding arbitrary bitstreams, it enables directly hiding and recovery of natural language messages, which benefits the steganographic techniques by expanding their expressive and potential use cases. Specifically, we formalize the Sentence-to-Image Steganography task, which aims to embed sentence-level (or even paragraph-level) natural language message into images such that the stego image remains visually indistinguishable from a benign image while allowing faithful recovery of the original message.
3.1 Problem Definition
The sentence-to-image steganography task aims to embed a secret message into a cover image to produce a stego image , from which can be fully recovered. This task emphasizes high capacity, seeking to conceal as much meaningful message as possible without degrading the visual quality of the cover image. Moreover, the secret messages can be any sentence-level, even paragraph-level data, encompassing more intricate expressions and longer text.
3.2 Benchmark
However, there is an absence of datasets specially designed for evaluating semantic steganography. To fill this gap, we construct the Invisible Text (IVT) dataset, which comprises a diverse set of sentence-level texts drawn from existing text datasets and generated by LLMs, serving as secret messages. We divide the messages into three levels of granularity: short messages (one sentence), medium messages (two to three sentences), and long messages (full articles), denoted as IVT-S, IVT-M, and IVT-L, respectively. Please refer to Appendix A for more details about the IVT dataset.
4 Our Method
4.1 Representation Information as Tokens
The representation of secret text is a critical aspect of semantic steganography. Previous methods convert each character of a word into bits using UTF encoding[tancik2020stegastamp, zhu2018hidden, wengrowski2019light]. However, these primitive approaches have two significant limitations. First, UTF encoding lacks semantic information, which is essential for representing sentence-level content. Second, representing a sentence as a bitstream is inefficient, as the resulting sequence becomes excessively long and difficult to process. We hypothesize that such primitive representations are particularly fragile when encoding richer semantic content. Therefore, it is necessary to explore more effective representations for sentence-level information.
To address these issues, we leverage the tokenizer of the LLM to convert the secret sentence into tokens. Tokenization is a widely used technique in NLP, where sentences are broken down into smaller units called Tokens. Tokens capture richer semantic information and facilitate modeling relationships between words. Moreover, tokenization significantly reduces sequence length, which benefits downstream processing. Using the tokenizer, we represent the secret sentence as a sequence of tokens. This approach not only preserves the semantic structure of the sentence but also ensures an efficient representation for embedding.
4.2 LLMs for Steganography
As shown in FigureΒ 2 the consists of three key modules: a pre-trained LLM for embedding and decoding the secret information, a Token-to-Patch MLP for aligning text features and image feature, and a Patch-to-Token MLP for embedding visual features to token embeddings. Notably, is a universal framework in which the LLM can be ANY pre-trained LLM.
Semantic embedding via LLMs. The LLM is recognized as the de facto standard architecture in the NLP community. Recent studies, however, have demonstrated that LLM can serve as a universal interface with different instructions in language to complete different tasks. In the computer vision community, researchers have explored integrating LLMs with vision backbones for segmentation, localization, and other visual tasks. Inspired by these works, we employ an LLM to generate embeddings conditioned by secret information, which will be inserted into the cover image to produce the stego image. Owing to the strong generalization ability of LLMs, we are able to compress the semantic information into a few embeddings.
Specifically, as illustrated in Figure 2, we design an embedding prompt and concatenate it with a secret message to instruct the LLM to embed the secret message into embedding, which we call secret message embeddings (SMEs). We call this strategy Prompt-Guide Embedding. The SMEs are then aligned with the image feature using the . After that, we reshape it and add to the cover image to get a stego image.
Universal insertion. To insert SMEs into the cover image, we use to align the token embeddings with the image features. Then, we adapt the UDH method [zhang2020udh], which directly adds the SMEs to the cover image to produce a stego image. Previous methods generate the stego image conditioned on both the cover image and the secret message. Empirically, however, the UDH method decouples the secret message from the carrier, resulting in smoother training and improved performance.
Decoding via LLMs. After receiving the stego image, we reshape it in patches and use to extract features and concatenate it with decoding prompt , which is used to instruct the LLM to decode the secret message. Finally, we get the recovery message from the LLM.
Template designing As illustrated in Figure 3, we carefully design the embedding and decoding prompts to instruct the LLM to embed and recover the secret message. We introduce two special tokens: and , which are used to separate secret messages from the prompt. We use the output embeddings of the LLM as SMEs. For clarity, we denote the location of the SMEs in the output of LLM as . During the decoding process, the extracted feature will be concatenated with the decode prompt as the input of LLM. We use as the placeholder for the feature extracted from stego image by .
4.3 Training of
Trainable parameters. To preserve the knowledge of the LLM, we use LoRA [hu2022lora] for fine-tuning. Additionally, because we introduce new tokens into the vocabulary (e.g. ), we also set the LLM token embeddings (), the LLM head (), and the projection layer trainable. By the way, the and are also trainable.
Stage 1: pre-training for steganography. As shown in FigureΒ 4, the training process consists of two stages. In the first stage, we train the LLM with LoRA, Token-to-Patch MLP , and Patch-to-Token MLP . In this stage, we DO NOT introduce the cover image. We wrap the secret message with and to construct the input sequence . Given the embedding prompt along with , we feed them into the LLM, which outputs a response token sequence , which can be formulated as:
We extract the last-layer embedding of the LLM as the secret embedding corresponding to the token in and apply to obtain , which aligns with the image patch dimension. Subsequently, will be processed by to align with the token embedding dimension of the LLM, yielding . To prevent the LLM from focusing solely on a subset of embeddings, we randomly mask with a mask ratio , which we refer to as the Mask Strategy. Finally, we concatenate it with decode prompt and feed it into the LLM to obtain the recovery message , which can be formulated as:
Stage 2: fine-tuning in stego image. In the second stage, we FREEZE the LLM and the , TRAIN the , and introduce the cover image . After obtaining , we reshape it from patch representation to the image representation , and add it to to obtain stego image . The entire process can be formulated as follow:
After receiving the stego image, we divide the it into patches to transform it into a 1D sequence and feed it into to obtain . Finally, we concatenate decode prompt with the and feed it into the LLM to get the recovery message , which can be formulated as follows:
Training objects. The is trained using the text generation loss and the embedding loss . The overall objective of each stage is defined as the weighted sum of these losses:
Specifically, is the auto-regressive cross-entropy (CE) loss for text generation, and is the L1 loss, which encourages the model to introduce minimal artifacts in the cover image. Those losses can be formulated as:
5 Experiments
5.1 Experimental Settings
Model variants. is a universal framework compatible with any pre-trained LLM. In this paper, we instantiate it with four models: Llama3.2-1B [grattafiori2024llama], Qwen2.5-0.5B [qwen2025qwen25technicalreport], Gemma3-1B [team2025gemma], and MiniCPM-1B [hu2024minicpm].
Implementation details. We employ AdamW optimizer [loshchilov2017decoupled]. The weights of the text generation loss and the embedding loss are set to 1.0 and 1.0, respectively. More details can be found on the Appendix B.1.
Datasets. For the secret message, we construct the training dataset using the CNN News Stories dataset [cnn_stories_dataset], news headlines [see2017get, ao-etal-2021-pens], and paper titles [farhangi2022protoformer]. For the cover image, we use the COCO training dataset [lin2014microsoft]. Please refer to Appendix B.2 for detailed information about training datasets. We evaluate the performance of steganographic models on the sentence-to-image steganography task using the IVT dataset. In experiments, we use the COCO test dataset as the cover image dataset. For more details about IVT, please refer to Appendix A.
Evaluation metrics. To measure the quality of the stego image, we use PSNR and SSIM to evaluate embedding performance. To assess recovery performance, we use word error rate (WER), BLEU-4 [papineni2002bleu], and ROUGE-L [lin2004rouge], and BERT-Score [zhang2019bertscore] to evaluate the semantic similarity between the ground truth and the recovery message.
Benchmarks. We compare with different methods, including DwtDct [Chen2009DigitalWA], StegaStamp [tancik2020stegastamp] and LanNet [lan2023robust]. We train StegaStamp and LanNet from scratch.
5.2 Sentence-to-Image Steganography Results
Sentence-to-Image results are shown in Table 1. Notably, while existing methods struggle with this task, our demonstrates strong performance.
Results of existing methods. Notably, existing methods are not suitable for the sentence-to-image steganography task, which fundamentally differs from watermarking in that it requires encoding a substantial amount of semantic information. As shown in TableΒ 1, StegaStamp (512-bit capacity) fails to recover secret messages in IVT-S, while DwtDct (768-bit capacity) performs well only on IVT-S but fails to recover messages in IVT-M, which exceed its capacity, not to mention handling IVT-L. However, LanNet (1024-bit capacity) also fails on IVT-L.
Results of . Existing methods fail to address this task effectively, whereas successfully handles secret messages containing complex and semantically rich natural language information.
On the IVT-S, all four variants demonstrate competitive performance compared to traditional baselines in secret recovery while significantly outperforming them in visual quality of stego image. Specifically, -MiniCPM-1B achieves the best overall performance, while -Qwen2.5-0.5B, -Llama3.2-1B, and -Gemma3-1B also deliver competitive results.
On the IVT-M, which involves longer and more complex message, -MiniCPM-1B maintains superior performance across most evaluation metrics. -Qwen2.5-0.5B and -Llama3.2-1B show slight performance degradation, while -Gemma3-1B demonstrates inferior semantic accuracy despite achieving high image quality.
On the IVT-L, all models exhibit performance degradation compared to their results on IVT-M. -Qwen2.5-0.5B get the top performance, -MiniCPM-1B and -Llama3.2-1B demonstrate moderate performance declines, and -Gemma3-1B exhibits the lowest semantic accuracy.
Experimental results demonstrate that all variants achieve strong performance on both IVT-S and IVT-M. However, for these models, IVT-L remains challenging, which comprises multi-sentence structures characterized by complex syntactic patterns and diverse lexical content. More details can be found in Appendix E.
| IVT-S | IVT-M | IVT-L | ||||||||||||||||
| Methods | Secret/Recovery | Cover/Stego | Secret/Recovery | Cover/Stego | Secret/Recovery | Cover/Stego | ||||||||||||
| WER | BLEU | ROUGE | BERT-S | PSNR | SSIM | WER | BLEU | ROUGE | BERT-S | PSNR | SSIM | WER | BLEU | ROUGE | BERT-S | PSNR | SSIM | |
| StegaStamp [tancik2020stegastamp] | 0.452 | 0.241 | 0.138 | 0.347 | 35.6 | 0.927 | - | - | - | - | - | - | - | - | - | - | - | - |
| DwtDct [chen2006dwt] | 0.007 | 0.699 | 0.996 | 0.996 | 31.8 | 0.898 | 0.701 | 0.490 | 0.713 | 0.829 | 27.7 | 0.776 | - | - | - | - | - | - |
| LanNet [lan2023robust] | 0.004 | 0.706 | 0.997 | 0.996 | 42.1 | 0.951 | 0.244 | 0.698 | 0.836 | 0.857 | 41.3 | 0.949 | 0.824 | 0.031 | 0.288 | 0.630 | 40.9 | 0.941 |
| S2LM-Qwen2.5-0.5B | 0.045 | 0.619 | 0.873 | 0.919 | 52.1 | 0.997 | 0.050 | 0.906 | 0.944 | 0.952 | 48.0 | 0.992 | 0.112 | 0.840 | 0.901 | 0.913 | 39.0 | 0.951 |
| S2LM-MiniCPM-1B | 0.046 | 0.635 | 0.923 | 0.959 | 59.8 | 0.999 | 0.061 | 0.913 | 0.962 | 0.965 | 52.6 | 0.996 | 0.197 | 0.762 | 0.859 | 0.889 | 46.9 | 0.988 |
| S2LM-Llama3.2-1B | 0.049 | 0.616 | 0.874 | 0.927 | 53.2 | 0.998 | 0.080 | 0.840 | 0.886 | 0.908 | 48.0 | 0.993 | 0.171 | 0.653 | 0.743 | 0.798 | 40.8 | 0.973 |
| S2LM-Gemma3-1B | 0.100 | 0.498 | 0.746 | 0.834 | 53.2 | 0.998 | 0.141 | 0.745 | 0.832 | 0.871 | 51.7 | 0.997 | 0.254 | 0.678 | 0.795 | 0.840 | 44.8 | 0.984 |
Qualitative results. As depicted in Figure 5, we provide a visualization of stego images and recovery results of -Qwen2.5-0.5B in different length of information. We mark the difference in red. As we can see, the recovery message and the ground truth keep high semantic similarity, while the stego image still has a good visual quality. More samples can be found in Appendix I.
| Capacity | Compress Ratio | Secret / Recovery | Cover / Stego | ||||
| WER | BLEU-4 | ROUGE-L | BERT-Score | PSNR | SSIM | ||
| 32 Token | 1:2 | 0.0331 | 0.9470 | 0.9752 | 0.9711 | 34.61 | 0.8759 |
| 64 Token | 1:1 | 0.0353 | 0.9523 | 0.9779 | 0.9694 | 31.64 | 0.7735 |
| 128 Token | 2:1 | 0.0855 | 0.8920 | 0.9476 | 0.9347 | 33.57 | 0.8536 |
| 256 Token | 4:1 | 0.4508 | 0.4847 | 0.7074 | 0.7454 | 29.53 | 0.5802 |
5.3 Capacity of the Cover Image
In this section, we conduct empirical experiments to investigate the capacity of the cover image under the framework. We analyze the capacity in terms of tokens, setting the secret message length to 32, 64, 128, and 256 tokens, while fixing the cover image resolution at . We train -Qwen2.5-0.5B from scratch for each setting. The cover image is divided into 64 patches, which is equal to the number of secret message embeddings (SMEs). For clearer analysis, we define the compression ratio as the number of SMEs divided by the number of patches.
| Two-Stage Training | Prompt | Embedding Loss | Mask Strategy | Secret Text / Recovery Text | Cover / Stego | ||||
| WER | BLEU-4 | ROUGE-L | BERT-Score | PSNR | SSIM | ||||
| 0.3162 | 0.5925 | 0.6781 | 0.6816 | 34.13 | 0.8156 | ||||
| β | 0.2210 | 0.6819 | 0.7403 | 0.8051 | 32.93 | 0.8460 | |||
[…]
| β | β | β | 0.2379 | 0.6703 | 0.7496 | 0.8119 | 34.24 | 0.7820 | |
[…]
As illustrated in TableΒ 2, the model maintains strong decoding performance at 32 and 64, 128 tokens, with BERT-Scores above 0.9. However, at 256 tokens, frequent decoding errors emerging. These results indicate that while the model has effective information compression ability, its capacity saturates beyond a certain threshold, leading to a noticeable degradation in recovery quality.
Based on these empirical findings, we propose a practical heuristic: each image patch can stably carry approximately two tokens without significant loss in decoding fidelity. This guideline provides an interpretable measure of capacity for future designs and scaling strategies, offering a concrete reference point for balancing payload length and image quality in sentence-to-image steganography.
5.4 Ablation Study
In this section, we conduct extensive ablation studies on -Qwen2.5-0.5B in the IVT-L to reveal the contribution of each component.
Effectiveness of two-stage training. As shown in Table 3, without two-stage training strategy, the struggles to recover the secret message. That may be because introducing the image modality at an early stage disrupts the LLMβs ability to learn steganographic mechanisms.
Effectiveness of the prompt. In this experiment, we do not use prompts to guide the LLM during the embedding and decoding processes. During embedding, we simply tokenize the secret message and input it into the LLM to obtain SMEs. Similarly, during decoding, we directly feed the features extracted by the Patch-to-Token MLP into the LLM to recover the original message. As shown in TableΒ 3, this approach leads to performance improvements of 7.3%, 7.5%, and 3.0% in BLEU-4, ROUGE-L, and BERT-Score, respectively. These results confirm that the prompt effectively guides the LLM in understanding and executing the sentence-to-image task.
Effectiveness of embedding loss. In the TableΒ 3, we show the contribution of the embedding loss . The is used to encourage the LLM to produce minimal artifacts in the cover image and is applied only in the first stage. However, experiment result shows that the introduction of the can improve the embedding quality with improvement of 15.61 dB and 0.165 in the PSNR and SSIM.
Effectiveness of mask strategy. In decoding process of the first stage, we mask some input tokens to encourage LLM distribute the information in the whole cover image. As shown in FigureΒ 6, the SMEs are evenly distributed after training with the mask strategy, and significant improvements in recovery quality are observed in TableΒ 3.
6 Conclusion
In this work, we propose a new steganographic task, sentence-to-image steganography, which is an instance of semantic steganography. Additionally, we present an evaluation benchmark, Invisible Text (IVT), comprising a diverse set of sentence-level texts drawn from existing datasets and generated by large language models (LLMs). Finally, we present , a novel steganographic framework that leverages LLMs to embed sentence-level information into images. We hope our work sheds new light on the advancement of semantic steganography.
Appendix A Invisible Text Dataset
To evaluate the effectiveness of semantic steganography, we construct a new benchmark dataset named InVisible Text (IVT). IVT is a large-scale dataset specifically designed for semantic steganography research, where the secret information consists of natural language text with varying degrees of complexity and length.
Specifically, we collect text information from various open-source datasets, including Translation, Question Answering, Sentiment Analysis, and Information Extraction et al. To facilitate systematic evaluation and controlled experimentation, we divide IVT into three subsets of increasing granularity: IVT-S, IVT-M, and IVT-L. These subsets differ in the length, structure, and semantic richness of the secret messages:
β’
IVT-S contains short phrases or word-level content targeting lightweight steganography scenarios.
β’
IVT-M contains complete sentences or short sentence pairs, balancing semantic richness and embedding complexity.
β’
IVT-L contains full natural-language paragraphs, serving as a high-capacity and semantically dense benchmark.
The word length distribution of IVT is shown in Figure 7. This three-granularity design enables comprehensive performance analysis across varying semantic loads, offering insights into the scalability, robustness, and text fidelity of semantic steganographic frameworks.
A.1 IVT-S
IVT-S is the simplest variant in IVT, focusing on short, low-capacity messages. Each secret message in IVT-S consists of a short phrase or a few words, typically ranging from 5 to 20 words. This setting reflects minimal semantic load, making it suitable for testing basic sentence-level embedding capabilities. IVT-S serves as a lightweight benchmark for evaluating models under low-information embedding scenarios.
The IVT-S dataset covers a wide range of short textual types, including real person names [name_country_origin_dataset], news headlines [ao-etal-2021-pens], paper titles [farhangi2022protoformer], Twitter comments [go2009twitter], and dialogue snippets [southparkdata]. For more details on IVT-S, please refer to Table 4.
| Category | Source |
License | ||||||||
| paper title | arxiv-10 [farhangi2022protoformer] | 9.7 | 592.2 | 4576 | 1000 | GPL-3.0 | ||||||||
| news title | PENS [ao-etal-2021-pens] | 10.3 | 506.9 | 5506 | 1000 | MSR License | ||||||||
| article title | examiner [examine_examiner_dataset] | 8.6 | 423.1 | 4873 | 1000 | CC0 1.0 | ||||||||
| people name | name&code [name_country_origin_dataset] | 2.1 | 113.8 | 857 | 1000 | - | ||||||||
| sentence | cmn_en [manythings_anki_dataset] | 6.2 | 247.3 | 1946 | 1000 | CC-BY 2.0 | ||||||||
| dialogue | South Park [southparkdata] | 4.5 | 202.9 | 1924 | 1000 | - | ||||||||
| comment | sentiment140 [go2009twitter] | 10.7 | 486.2 | 4374 | 1000 | - | ||||||||
| overall | - | 7.4 | 367.5 | 19941 | 7000 | - |
A.2 IVT-M
IVT-M is a medium-granularity subset in IVT, positioned between the short messages of IVT-S and the full-paragraph texts of IVT-L. Each secret message in IVT-M consists of coherent sentences with a length of about 50 to 100 words. This setting introduces moderate semantic complexity and content richness, making it ideal for evaluating models under typical sentence-level steganography scenarios. IVT-M reflects more realistic use cases than IVT-S by requiring the model to encode and recover messages with non-trivial structure and meaning.
Messages in IVT-M are sourced from a variety of domains, including CNN news abstracts [see-etal-2017-get], Wikipedia summaries [wikidump], IMDB reviews [imdb_reviews_dataset]. For more details about IVT-M, please refer to Table 5.
| Category | Source |
License | ||||||||
| paper title | arxiv-10 [farhangi2022protoformer] | 22.0 | 1210.0 | 7776 | 1000 | GPL-3.0 | ||||||||
| new title | PENS [ao-etal-2021-pens] | 23.0 | 1369.4 | 8003 | 1000 | MSR License | ||||||||
| news highlight | CNN and Daily mail [see2017get] | 50.3 | 2303.6 | 13119 | 1000 | Apache 2.0 | ||||||||
| translate text | WMT-14 [bojar-EtAl:2014:W14-33] | 33.8 | 1610.1 | 6760 | 1000 | - | ||||||||
| joke story | stupid joke [pungas] | 47.6 | 2042.1 | 10637 | 1000 | - | ||||||||
| translate text | europarl [koehn2005europarl] | 34.0 | 1618.8 | 6735 | 1000 | - | ||||||||
| dialogue | South Park [southparkdata] | 31.7 | 1367.1 | 3896 | 1000 | - | ||||||||
| comment | hate text [hateoffensive] | 23.4 | 989.9 | 6564 | 1000 | MIT | ||||||||
| moral | moral story [guan2022corpus] | 32.5 | 1369.4 | 3179 | 1000 | CC BY-NC-SA 3.0 | ||||||||
| overall | - | 33.1 | 1542.3 | 42923 | 9000 | - |
A.3 IVT-L
IVT-L represents the largest and most challenging subset of the IVT dataset, designed to evaluate a modelβs ability to embed and recover long-form semantic content. Each secret message in IVT-L consists of a full paragraph containing multiple sentences with coherent structure, complex syntax, and a diverse vocabulary. Compared to IVT-S and IVT-M, IVT-L significantly enhances the semantic richness and information density of the hidden text, thereby imposing greater demands on a modelβs capacity, fidelity, and ability to preserve semantics. This subset serves as a benchmark for evaluating whether steganographic models can scale to real-world scenarios that involve concealing full-length, meaningful textual content.
IVT-L includes long-form messages drawn from a variety of sources, including narrative news passages, short stories [guan2022corpus, eldan2023tinystories], research paper abstracts [farhangi2022protoformer], IMDB reviews [imdb_reviews_dataset]. These samples span multiple domains and writing styles, providing a comprehensive benchmark for evaluating sentence-to-image steganography in high-capacity settings. For more details about IVT-L, please refer to Table 6.
| Category | Source |
License | ||||||||
| paper abstract | arxiv-10 [farhangi2022protoformer] | 156.55 | 8512.75 | 22550 | 1000 | GPL-3.0 | ||||||||
| wiki abstract | wiki [wikidump] | 140.67 | 6962.01 | 25117 | 1000 | GFDL/CC BY-SA | ||||||||
| story | stupid joke [pungas] | 153.37 | 6725.4 | 23701 | 1000 | - | ||||||||
| story | moral story [guan2022corpus] | 225.92 | 9679.95 | 18412 | 1000 | - | ||||||||
| story | Tiny Story [eldan2023tinystories] | 172.33 | 7073.81 | 10477 | 1000 | - | ||||||||
| user review | IMDB [imdb_reviews_dataset] | 194.06 | 8761.54 | 30569 | 1000 | IMDB License | ||||||||
| user review | yelp [asghar2016yelp] | 128.16 | 5542.79 | 17244 | 1000 | Yelp License | ||||||||
| essay | asap-aes [asap_aes_dataset] | 199.43 | 8734.33 | 15142 | 1000 | - | ||||||||
| overall | - | 171.3113 | 7749.073 | 111662 | 8000 | - |
Appendix B Experimental Details
To explore the feasibility and performance of sentence-to-image steganography powered by large language models, we propose three variants of our framework, namely -Qwen2.5-0.5B, -Llama3.2-1B, and -MiniCPM-1B. These models are built upon different backbone LLMs with varying scales and capabilities, providing diverse perspectives on model behavior across capacity and architecture. In the following sections, we detail the training procedures and dataset construction strategies used for each model, including how the secret messages are formatted, encoded, and paired with cover images to form effective training samples. This enables a thorough evaluation of how different LLMs perform under the semantic steganography task.
B.1 Implementation Details
LoRA Config. We employ Low-Rank Adaptation (LoRA) with a rank to fine-tune the pretrained LLM using the peft library, injecting trainable low-rank matrices into the query and key projection layers (q_proj, k_proj). The LoRA scaling factor is set to 32 with a dropout rate of 0.1.
Trainer. We train the model using the Trainer of the transformers library. The is optimized via the AdamW optimizer with a learning rate of 2e-4 and weight decay of 0.01 for regularization. The learning rate follows a cosine scheduler with 500 warmup steps to stabilize early training. We use mixed-precision training to accelerate convergence while maintaining numerical stability. The two-stage training shares the same config.
Computation Platform. All experiments are conducted on NVIDIA A800 80GB GPUs with CUDA Version 12.4.
-Qwen2.5-0.5B. -Qwen2.5-0.5B is built upon the Qwen2.5-0.5B backbone [qwen2025qwen25technicalreport] and fine-tuned using the LoRA config mentioned above. The training strategy is detailed in Section 4.3. In Stage 1, we use a batch size of 14 and train for one epoch. In Stage 2, we train for 6,000 iterations with a batch size of 14.
-Llama3.2-1B. -Llama3.2-1B is built upon the Llama3.2-1B backbone [grattafiori2024llama] and fine-tuned using the LoRA config mentioned above. The training strategy is detailed in Section 4.3. In Stage 1, we use a batch size of 14 and train for one epoch. In Stage 2, we train for 6,000 iterations with a batch size of 14.
-MiniCPM-1B. -MiniCPM-1B is built upon the MiniCPM-1B backbone [hu2024minicpm] and fine-tuned using the LoRA config mentioned above. The training strategy is detailed in Section 4.3. In Stage 1, we use a batch size of 24 and train for one epoch. In Stage 2, we train for 5,000 iterations with a batch size of 24.
-Gemma3-1B. -Gemma3-1B is built upon the Gemma3-1B backbone [team2025gemma] and fine-tuned using the LoRA config mentioned above. The training strategy is detailed in Section 4.3. In Stage 1, we use a batch size of 14 and train for one epoch. In Stage 2, we train for 6,000 iterations with a batch size of 14.
StegaStamp. We retrain StegaStamp [tancik2020stegastamp] with extended capacity (from 100 to 512 bits) on cover images with 256Γ256 resolution using the Flickr2K dataset. The model is optimized with AdamW with learning rate set to 2e-4 for 15,000 iterations with batch size set to 16. While achieving stable convergence, we observe fundamental limitations: the 512-bit encoding yields only ~70% bit accuracy, resulting terrible performance in IVT. This suggests current spatial-domain watermarking architectures struggle with high-capacity payloads with primitive representation of secret message such as UTF codes.
LanNet. We retrain LanNet [lan2023robust] on cover images with 256Γ256 resolution using the DIV2K training dataset. The model is optimized with AdamW with learning rate set to 2e-4 for 2,000 epochs with batch size set to 16.
B.2 Details about Training Dataset
To construct the training dataset for the aforementioned models, we curated a diverse collection of textual data from multiple sources. Specifically, we selected 14,000 highlights from the CNN/DailyMail dataset [see-etal-2017-get], 5,000 news headlines from the PENS dataset [ao-etal-2021-pens], 5,000 scientific titles from the ArXiv-10 dataset [farhangi2022protoformer], and 4,000 story paragraphs from the CNN Story corpus [cnn_stories_dataset]. This composition ensures a broad range of semantic structures and topics, facilitating robust model training across different text granularities.
Importantly, we ensure that there is no overlap between the training set and the proposed IVT benchmark datasets, thereby preventing data leakage and preserving the integrity of the evaluation. For more details about IVT-L, please refer to Table 7.
| Category | Source |
| news headlines | PENS [ao-etal-2021-pens] | 21.6 | 1296.3 | 5000 | ||||||
| paper title | arxiv-10 [farhangi2022protoformer] | 25.1 | 1402.0 | 5000 | ||||||
| news highline | CNN/DailyMail [see-etal-2017-get] | 55.8 | 2826.4 | 14000 | ||||||
| new story | CNN Story [cnn_stories_dataset] | 362.6 | 11024.6 | 4000 |
Appendix C Steganalysis
C.1 Statistical Steganalysis
We use StegExpose [Boehm2014StegExposeA] to measure the anti-steganalysis ability. The receiver operating characteristic (ROC) curve in FigureΒ 8 shows that has high security and can fool the StegExpose.
Appendix D VLMs in
In the previous experiments, all models were based on pure text-only large language models. To further explore the applicability of vision-language models (VLMs) to the sentence-to-image steganography task, we conducted additional experiments using Qwen2-vl [Qwen2-VL] as a representative VLM. However, the results showed that Qwen2-vl failed to fit the training data.
We hypothesize that this underperformance stems from the fundamental difference between visual understanding and steganography. While VLMs are primarily optimized for visual tasks, which focusing on understanding objects, scenes, and spatial relationships within images. However, steganography emphasizes decoding secret messages by focusing on subtle, imperceptible signals embedded within images. Unlike typical vision-language alignment tasks, steganography does not require high-level visual semantics but rather the capacity to detect and interpret fine-grained patterns that are deliberately encoded. This highlights the need for models specifically tailored to the semantic embedding and recovery processes intrinsic to image-based steganography.
In experiments, the -Qwen2-vl-2B model can output the and tokens, but the recovery information is totally different from the ground truth. That may because the knowledge in the vision encoder seriously hindering the language model to get exact secret information.
Appendix E Experiments on Generative Message
To further assess the ability of our , we construct a new dataset named IVTG, which is entirely synthesized by the LLM. Specifically, we leverage the DeepSeek [deepseek2024] API to generate the textual content, as illustrated in Appendix Section F. Similar to the original IVT dataset, we divide IVTG into three subsets based on the length of the secret text.
As shown in Table 8, our model achieves comparable performance on IVTG to its performance on the original IVT dataset. This consistency demonstrates the modelβs ability to generalize to data generated by large language models, further validating the effectiveness of our approach.
Appendix F Codes for Collecting IVTG
| IVTG-S | IVTG-M | IVTG-L | ||||||||||||||||
| Methods | Secret/Recovery | Cover/Stego | Secret/Recovery | Cover/Stego | Secret/Recovery | Cover/Stego | ||||||||||||
| WER | BLEU | ROUGE | BERT-S | PSNR | SSIM | WER | BLEU | ROUGE | BERT-S | PSNR | SSIM | WER | BLEU | ROUGE | BERT-S | PSNR | SSIM | |
[…]
Appendix G Limitations
While demonstrates strong performance in encoding and recovering semantically rich text, it exhibits limitations when handling non-semantic or low-semantic content, such as random digit strings, binary sequences, or nonsensical phrases.
This limitation contrasts with some watermarking-based frameworks, which are designed to handle raw bit-level information and therefore remain agnostic to the semantic nature of the payload. In contrast, leverages a language modeling backbone that inherently assumes the presence of syntactic and semantic structures in the secret message.
Appendix H Broader Impacts
Semantic steganography introduces a novel paradigm that leverages carriers to conceal the structure and semantics of natural language, enabling new possibilities beyond traditional bit-level encoding schemes. Our method, , focuses on embedding textual payloads within images. However, the underlying principle is extensible to a wide range of modalities.
Specifically, this framework can be generalized to video, audio, or even 3D environments. Embedding meaningful language into other media formats presents opportunities for developing cross-modal steganographic systems capable of high-capacity, human-aligned, and content-aware data hiding.
We hope that this work not only advances semantic steganography from a technical perspective but also encourages discussion around its responsible use, ethics, and governance frameworks in future applications.
Appendix I Additional Results
NeurIPS Paper Checklist
1.
Claims
Question: Do the main claims made in the abstract and introduction accurately reflect the paperβs contributions and scope?
Answer: [Yes]
Justification: The abstract and introduction clearly state the proposed concept of semantic steganography, introduce the sentence-to-image task, and present the , accurately reflecting the paperβs contributions and scope.
Guidelines:
β’
The answer NA means that the abstract and introduction do not include the claims made in the paper.
β’
The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
β’
The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
β’
It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
β’
2.
Limitations
Question: Does the paper discuss the limitations of the work performed by the authors?
Answer: [Yes]
Justification: We discusses its limitations in the appendix G, acknowledging areas for future improvement and potential constraints of the proposed approach.
Guidelines:
β’
The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
β’
The authors are encouraged to create a separate "Limitations" section in their paper.
β’
The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
β’
The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
β’
The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
β’
The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
β’
If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
β’
While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that arenβt acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
β’
3.
Theory assumptions and proofs
Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
Answer: [No]
Justification: No theoretical result
Guidelines:
β’
The answer NA means that the paper does not include theoretical results.
β’
All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
β’
All assumptions should be clearly stated or referenced in the statement of any theorems.
β’
The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
β’
Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
β’
Theorems and Lemmas that the proof relies upon should be properly referenced.
β’
4.
Experimental result reproducibility
Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
Answer: [Yes]
Guidelines:
β’
The answer NA means that the paper does not include experiments.
β’
If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
β’
If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
β’
Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
β’
While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
β’
5.
Open access to data and code
Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
Answer: [Yes]
Justification: The IVT dataset is included in supplemental material, the code of will be released after the paper being accepted.
Guidelines:
β’
The answer NA means that paper does not include experiments requiring code.
β’
Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
β’
While we encourage the release of code and data, we understand that this might not be possible, so βNoβ is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
β’
The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
β’
The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
β’
The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
β’
At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
β’
Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
β’
6.
Experimental setting/details
Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
Answer: [Yes]
Guidelines:
β’
The answer NA means that the paper does not include experiments.
β’
The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
β’
The full details can be provided either with the code, in appendix, or as supplemental material.
β’
7.
Experiment statistical significance
Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
Answer: [No]
Justification: We do not report error bars following prior arts [jing2021hinet, zhu2018hidden, ke2024stegformer, baluja2017hiding, lu2021large].
Guidelines:
β’
The answer NA means that the paper does not include experiments.
β’
The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
β’
The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
β’
The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
β’
The assumptions made should be given (e.g., Normally distributed errors).
β’
It should be clear whether the error bar is the standard deviation or the standard error of the mean.
β’
It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
β’
For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
β’
If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
β’
8.
Experiments compute resources
Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
Answer: [Yes]
Justification: The training costs are reported in Appendix Section B.1.
Guidelines:
β’
The answer NA means that the paper does not include experiments.
β’
The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
β’
The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
β’
The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didnβt make it into the paper).
β’
9.
Code of ethics
Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
Answer: [Yes]
Justification: The research in the paper conform with the NeurIPS Code of Ethics.
Guidelines:
β’
The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
β’
If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
β’
The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
β’
10.
Broader impacts
Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
Answer: [Yes]
Justification: The broader impacts have been discussed in Appendix Section H.
Guidelines:
β’
The answer NA means that there is no societal impact of the work performed.
β’
If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
β’
Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
β’
The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
β’
The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
β’
If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
β’
11.
Safeguards
Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
Answer: [N/A]
Justification: The paper poses no such risks.
Guidelines:
β’
The answer NA means that the paper poses no such risks.
β’
Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
β’
Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
β’
We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
β’
12.
Licenses for existing assets
Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
Answer: [Yes]
Justification: Please refer to Appendix Section A.
Guidelines:
β’
The answer NA means that the paper does not use existing assets.
β’
The authors should cite the original paper that produced the code package or dataset.
β’
The authors should state which version of the asset is used and, if possible, include a URL.
β’
The name of the license (e.g., CC-BY 4.0) should be included for each asset.
β’
For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
β’
If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
β’
For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
β’
If this information is not available online, the authors are encouraged to reach out to the assetβs creators.
β’
13.
New assets
Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
Answer: [N/A]
Justification: This paper does not release new assets.
Guidelines:
β’
The answer NA means that the paper does not release new assets.
β’
Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
β’
The paper should discuss whether and how consent was obtained from people whose asset is used.
β’
At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
β’
14.
Crowdsourcing and research with human subjects
Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
Answer: [N/A]
Justification: This paper does not involve crowdsourcing nor research with human subjects.
Guidelines:
β’
The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
β’
Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
β’
According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
β’
15.
Institutional review board (IRB) approvals or equivalent for research with human subjects
Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
Answer: [N/A]
Justification: This paper does not involve crowdsourcing nor research with human subjects.
Guidelines:
β’
The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
β’
Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
β’
We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
β’
For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
β’
16.
Declaration of LLM usage
Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
Answer: [Yes]
Guidelines:
β’
The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
β’
Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.
β’

### v2

@@
 Abstract
-Although steganography has made significant advancements in recent years, it still struggles to embed semantically rich, sentence-level information into carriers.
-However, in the era of AIGC, the capacity of steganography is more critical than ever.
-In this work, we present Sentence-to-Image Steganography, an instance of Semantic Steganography, a novel task that enables the hiding of arbitrary sentence-level messages within a cover image.
-Furthermore, we establish a benchmark named Invisible Text (IVT), comprising a diverse set of sentence-level texts as secret messages for evaluation.
-Finally, we present : Semantic Steganographic Language Model, which utilizes large language models (LLMs) to embed high-level textual informationβsuch as sentences or even paragraphsβinto images.
-Unlike traditional bit-level counterparts, enables the integration of semantically rich content through a newly designed pipeline in which the LLM is involved throughout the entire process.
-Both quantitative and qualitative experiments demonstrate that our method effectively unlocks new semantic steganographic capabilities for LLMs.
-The source code will be released soon.
+Despite remarkable progress in steganography, embedding semantically rich, sentence-level information into carriers remains a challenging problem.
+In this work, we present a novel concept of Semantic Steganography, which aims to hide semantically meaningful and structured content, such as sentences or paragraphs, in cover media.
+Based on this concept, we present Sentence-to-Image Steganography as an instance that enables the hiding of arbitrary sentence-level messages within a cover image.
+To accomplish this feat, we propose : Semantic Steganographic Language Model, which leverages large language models (LLMs) to embed high-level textual information into images.
+Unlike traditional bit-level approaches, redesigns the entire pipeline, involving the LLM throughout the process to enable the hiding and recovery of arbitrary sentences.
+Furthermore, we establish a benchmark named Invisible Text (IVT), comprising a diverse set of sentence-level texts as secret messages to evaluate semantic steganography methods.
+Experimental results demonstrate that effectively enables direct sentence recovery beyond bit-level steganography.
+The source code and IVT dataset will be released soon.
 1 Introduction
-In an age long past, under the shadow of the Persian Empire, Histiaeus, ruler of Miletus, devised a plan that would etch his name in the annals of history.
-Under the watch of the imperial, he tattooed a secret message onto the scalp of a loyal slave, waiting for the hair to regrow and conceal the message beneath a veil of innocence.
-When the time was right, the slave crossed enemy lines, carrying the message not in his hands but under his skin.
-The message sparked a revolution, a call to arms that would ignite the Ionian Revolt.
-This tale, immortalized by Herodotus in the Histories, remains one of the earliest recorded feats of Steganography - the art of hiding information in plain sight.
-Fast forward today, on the verge of entering the Artificial Intelligence Generated Content (AIGC) era, the principles of steganography have never been more relevant, with the prosperity of various generative models (e.g., Stable Diffusion [rombach2021highresolution], ChatGPT [openai_chatgpt], and Llama [touvron2023llama]), which can produce highly realistic text [bai2023qwentechnicalreport, qwen2025qwen25technicalreport], images [rombach2021highresolution, ramesh2022hierarchicaltextconditionalimagegeneration] and videos [liu2024sorareviewbackgroundtechnology], and even execute tasks with human-like proficiency [durante2024agentaisurveyinghorizons].
-Their widespread application has blurred the line between fact and forgery, raising numerous security and legal concerns.
-The application of the generative model may foster the generation of biased content [ling2025bias, serouis2024exploring, kumar2025no], image manipulation [qu2024towards], and the creation of fake news [wu2024fake].
-Moreover, the human-like content generated by those models is flooding the entire Internet, seriously hindering the determination of the credibility of evidence in forensics [yin2025digital].
-Therefore, there is an urgent need for effective measures to prevent the large-scale misuse of generative models.
-To address this problem, embedding forensics information into the generated content by steganography could be the most effective method.
-As the real-world application of steganography, embeds secret information that is typically difficult to detect and remove but can be easily retrieved [hartung2002multimedia].
-With this technique, it is possible to determine whether the generative model creates the content quickly.
-Moreover, the extracted information can also be used for copyright identification and source tracing, thereby protecting intellectual property rights and helping locate the infringer.
-Nowadays, various strategies have been developed for text [pan2024markllm, zhang2024remark, liang2024watermarkingtechniqueslargelanguage], images [wen2023tree, fernandez2023stable], video [lin2024automatic, luo2023dvmark], and audio [zhang2024v2a] generated by generative models.
-As a primary medium of visual communication, images are widely shared on the Internet, promoting the development of digital image steganography.
-Over years of research, numerous methods have been proposed[zhu2018hidden, wengrowski2019light, tancik2020stegastamp, wen2023tree, cox2007digital, wang2001image, patel2012steganography].
-However, most existing approaches focus on embedding low-level binary or character-based information, neglecting the semantic relationships inherent in natural language.
-Consequently, these methods exhibit limited capacity and poor expressivity, preventing them from embedding sentence-level messages.
-Those methods underutilize the representational potential of modern high-resolution images and fail to meet the increasing capacity demands of watermarking in the era of generative AI.
-For further analysis, we divide the current steganography [zhu2018hidden, wengrowski2019light, tancik2020stegastamp, patel2012steganography] pipeline into three processes: the encoding process, which transforms the secret information into a suitable representation; the insertion process, which embeds the message into the cover image to produce a stego image; and the decoding process, which extracts the secret message from the stego image.
-As illustrated in Figure 1, three significant challenges hinder current steganographic methods from embedding substantial semantic information into images: 1) In the encoding process, secret messages are typically converted into bitstreams based on ASCII or UTF formats.
-This method is simple but primitive, neglects the semantic information between words, has a low information density, and is unsuitable for encoding large information such as sentences; 2) In the insertion process, the encoder takes the secret message and cover image as input to generate a stego image.
+Steganography is the art of hiding information in plain sight.
+Historically, it was used as a means of covert communication during wartime.
+Nowadays, these techniques have found broader applications in copyright protection [80], tamper detection [79, 81], and secure communication [72].
+With the advent of deep learning, steganography has achieved remarkable progress in recent years.
+However, despite this apparent prosperity, existing steganographic models can hide only few words of secret information.
+This limitation primarily arises from their design paradigm: the secret message must first be converted into a bitstream before being processed by the model.
+Consequently, current methods are fundamentally designed to embed binary data [68, 60, 9], which severely restricts their capacity and leads to low communication efficiency in practical applications.
+For further analysis, we take image steganography as a case study and divide the steganographic pipeline into three processes, as illustrated in Fig.Β 1.
+There are three significant limitations that hinder existing methods from embedding larger amounts of information into images: 1) In the encoding process, messages are converted into bitstreams via ASCII or UTF encoding.
+This method is simple but primitive, neglects the semantic relationships between words, has a low information density, and is unsuitable for encoding large amounts of information, such as sentences.
+2) In the insertion process, the encoder takes the secret message and cover image as input to generate a stego image.
 The interaction between two fundamentally different modalities (bitstreams and images) results in suboptimal insertion performance.
-3) In the decoding process, due to the aforementioned limitations, the decoder can only recover bitstreams, restricting the hidden information to the word level.
-The three significant challenges mentioned above hinder the existing steganography to increase the secret message capacity to include semantic information, such as complete sentences.
-To address these challenges, we introduce the broader concept of Semantic Steganography, which focuses on embedding semantically meaningful and structured contentβsuch as sentences or paragraphsβinto carriers.
+3) In the decoding process, due to the aforementioned constraints, the decoder can only recover bitstreams, restricting the hidden information to the word level.
+These practical limitations call for a rethinking of steganography: Are steganographic frameworks inherently limited to hiding low-level bitstream information?
+Motivated by this question, we introduce the broader concept of Semantic Steganography, which focuses on embedding semantically meaningful and structured content, such as natural language, into carriers.
 Building on this idea, we formalize the Sentence-to-Image Steganography task as an instance of semantic steganography.
-This task requires hiding sentence-level or even paragraph-level semantic information into a cover image.
+This task requires hiding sentence-level or even paragraph-level information in a cover image.
 Notably, the secret message can consist of arbitrary textual content of any length.
-To accomplish this task, the model must possess two key capabilities: 1) encoding large and complex information into an appropriate representation and 2) effectively handling two fundamentally different modalities (text and image).
-Recent advancements in the development of large language models (LLMs) have spurred efforts to integrate LLMs with vision backbones, leading to the rise of vision-language models (VLMs) [zhang2024vision, ye2024mplug, li2022mplug, liu2024improved, liu2023visual].
-However, most of these models focus on vision understanding and text generation.
-This raises an important question: Can LLMs be used directly as a universal steganographic tool to embed sentence-level information into carriers such as images?
-To this end, we propose the Semantic Steganographic Language Model (), an innovative steganographic framework.
-To the best of our knowledge, this is the first work to explore hiding sentence-level and even paragraph-level information in images.
-Specifically, we redesign the entire pipeline used in current watermarking strategies.
-As illustrated in FigureΒ 1, we first encode the secret sentence into tokens using a tokenizer.
-Second, during the insertion process, we use the LLM to generate secret embeddings without relying on image priors, thereby eliminating interference between the cover image and the secret message.
-Finally, we use the same LLM to decode the secret message from the stego image.
-In our empirical experiments, can conceal up to 500 words of text into a single image with a resolution of , breaking the boundary between image and text.
-Additionally, we evaluate the steganographic performance of different LLMs within the framework and verify its versatility across various language models.
+To accomplish this task, the model must possess two key capabilities: encoding large and complex information into an appropriate representation and effectively handling two fundamentally different modalities.
+Recent advancements in large language models (LLMs) have spurred research on integrating them with vision backbones, leading to the rise of vision-language models (VLMs) [76, 70, 37, 42, 43].
+These studies demonstrate that LLMs exhibit strong capabilities in semantic understanding and cross-modal reasoning, which precisely align with the requirements of semantic steganography.
+This alignment highlights the potential of LLMs to serve as a foundation for developing a universal steganographic framework that embeds high-level semantics within carriers.
+Building on this insight, we propose the Semantic Steganographic Language Model (), a novel framework that leverages the semantic understanding capability of LLMs for steganography.
+To address the three challenges mentioned above, completely redesigns the entire pipeline of current methods, as illustrated in Fig.Β 1.
+1) In the encoding process, we first transform the secret textual information into a sequence of discrete tokens using a tokenizer.
+This process not only captures the semantic content of the sentence but also preserves the relationships between words, enabling more effective and structured representation of the secret message.
+2) During the insertion process, we leverage LLM to generate secret embeddings without relying on any prior of the cover image.
+By decoupling secret embeddings from the carrier, prevents potential interference between the hidden message and the cover image.
+3) In the decoding process, we employ the same LLM to decode the secret message from the stego image.
+Overall, provides a unified framework to handle the sentence-to-image task, demonstrating strong capability in encoding and recovering semantically rich messages without compromising visual integrity.
+In our qualitative experiments, successfully conceals up to 500 words of text into a single image with a resolution of , breaking the boundary between image and text.
+We further evaluate the steganographic performance of different LLMs within the framework and verify its versatility across various models.
 Additionally, we establish a benchmark for semantic steganography called Invisible Text (IVT), which we use to validate the effectiveness of the sentence-to-image framework in this paper.
@@
 β’
-We propose a new steganographic taskβSentence-to-Image Steganography, which requires embedding sentence-level information into a cover image.
+We propose the concept of Semantic Steganography and formalize Sentence-to-Image Steganography as an instance, redefining the traditional goal of steganography from embedding low-level data to concealing meaningful language content within cover media.
 β’
@@
 β’
-We establish a benchmark for the semantic steganography named Invisible Text (IVT), containing thousands of sentence-level samples from diverse sources.
+We establish a benchmark for semantic steganography named Invisible Text (IVT), containing thousands of samples from diverse sources.
 This benchmark is essential for evaluation and encourages the community to further explore the potential of semantic steganography.
-2 Related Works
+2 Related Work
 2.1 Image Steganography
 Steganography is a technique for concealing secret information within a carrier while preserving its visual appearance and avoiding detection.
-Traditional image steganography methods, such as the Least Significant Bit (LSB) algorithm [wang2001image] and frequency-domain approaches [patel2012steganography, chen2006dwt] (e.g., DCT and DWT), embed secret data by modifying the pixel values or transform coefficients of the cover image.
-However, these methods have limited capacity and struggle to meet the requirements of modern steganography, whereas deep learning-based methods perform better.
-Recent advances in deep learning have produced more powerful steganographic techniques, such as HiDDeN [zhu2018hidden], LFM [wengrowski2019light], and StegaStamp [tancik2020stegastamp], which can embed bit-level information in images with robustness against various manipulations.
-Moreover, methods like HiNet [jing2021hinet], ISN [lu2021large], LanNet [lan2023robust], and StegFormer [ke2024stegformer] offer greater capacity, enabling the concealment of an image in carrier with the same resolution.
+Traditional image steganography methods, such as the Least Significant Bit (LSB) algorithm [65] and frequency-domain approaches [48, 12], embed secret data by modifying the pixel values or transform coefficients of the cover image.
+However, these hand-crafted methods tend to introduce statistical artifacts that can be easily detected by steganalysis, whereas deep learning-based methods achieve higher security.
+Recent advances in deep learning have produced more powerful steganographic techniques, such as HiDDeN [83], LFM [68], and StegaStamp [60], which can embed bit-level information in images with robustness against various manipulations.
+Moreover, methods like HiNet [29], ISN [45], LanNet [36], and StegFormer [31] offer greater capacity, enabling the concealment of an image in a carrier with the same resolution.
+Recently, several works have explored utilizing diffusion models for steganography, such as DGADM-GIS [73], RoSteALS [9], and Latent Watermark [46].
 Despite these advancements, existing methods primarily focus on embedding secret information at a low level (bits or images).
@@
 2.2 Large Language Models in Computer Vision
-The prosperity of the current large language model (LLM) has pushed the whole community of the Natural Language Processing (NLP) into new boundaries.
-The computer vision community has also noticed the tremendous success of the practice of LLM.
+The prosperity of current LLMs has pushed the entire Natural Language Processing community into new boundaries.
+The computer vision community has also recognized the tremendous success of LLMs.
 Consequently, lots of the latest research focuses on exploring the integrity of the LLM with the vision backbone.
-BLIP [li2022blip] proposes a Multi-modal mixture of Encoder-Decoder (MED), while BLIP2 [li2023blip] introduces a frozen LLM to bootstrap vision-to-language learning, and proposes a Q-Former to bridge the modality gap.
-LLaVA [liu2023visual] connects the vision backbone and Llama [touvron2023llama] and tunes the LLM with multi-modal language-image instruction-following data, achieving the general-purpose visual and language understanding.
+BLIP [39] proposes a Multi-modal mixture of Encoder-Decoder (MED), while BLIP2 [38] introduces a frozen LLM to bootstrap vision-to-language learning, and proposes a Q-Former to bridge the modality gap.
+LLaVA [43] connects the vision backbone and Llama [62] and tunes the LLM with multi-modal language-image instruction-following data, achieving the general-purpose visual and language understanding.
 The methods above explore the LLM application in visual understanding, often seen as a high-level vision task.
 However, the LLM can also enhance the performance of the low-level vision backbone.
-LISA [Lai2023LISARS] uses a novel segmentation token and unlocks the segmentation capability of the LLM, which can be used for reasoning segmentation tasks.
+LISA [35] uses a novel segmentation token and unlocks the segmentation capability of the LLM, which can be used for reasoning segmentation tasks.
 Xu et al.
-used localized image captioning data to pretrain PixelLLM to equip it with localization capability [xu2024pixel].
+used localized image captioning data to pretrain PixelLLM to equip it with localization capability [69].
 Zheng et al.
-froze LLM to solve a range of low-level vision tasks, demonstrating the LLMβs strong generalization capability [zheng2024lm4lv].
-In this paper, we investigate the steganographic capabilities of LLMs, which has not been explored in the previous works.
-Experiments show that LLMs can serve as a steganographic tool to embed sentence-level information into a cover image.
-3 Sentence-to-Image Steganography
+froze LLM to solve a range of low-level vision tasks, demonstrating the LLMβs strong generalization capability [82].
+While several studies have attempted to apply LLMs to linguistic steganography [66, 5, 54], most of them rely on manually designed sampling strategies.
+In this paper, we integrate LLM into the entire steganographic pipeline, directly utilizing LLM to generate secret embeddings directly, and successfully scale the steganographic capacity from bit to sentence-level.
+3 Semantic Steganography
 The whole steganographic community has traditionally focused on low-level information hiding (e.g., at the bit or image level), overlooking the semantic structure of the hidden information.
@@
 This paradigm redefines what kind of information can be hidden.
-Rather than merely embedding arbitrary bitstreams, it enables directly hiding and recovery of natural language messages, which benefits the steganographic techniques by expanding their expressive and potential use cases.
-Specifically, we formalize the Sentence-to-Image Steganography task, which aims to embed sentence-level (or even paragraph-level) natural language message into images such that the stego image remains visually indistinguishable from a benign image while allowing faithful recovery of the original message.
+Rather than merely embedding arbitrary bitstreams, it enables the direct hiding and recovery of natural language messages, which benefits the steganographic techniques by expanding their expressive and potential use cases.
 3.1 Problem Definition
-The sentence-to-image steganography task aims to embed a secret message into a cover image to produce a stego image , from which can be fully recovered.
-This task emphasizes high capacity, seeking to conceal as much meaningful message as possible without degrading the visual quality of the cover image.
-Moreover, the secret messages can be any sentence-level, even paragraph-level data, encompassing more intricate expressions and longer text.
+We define semantic steganography as the task of hiding semantic content within a carrier while keeping it undetectable.
+In this work, the semantic content of interest is expressed in natural language, specifically as complete sentences that convey coherent and interpretable meaning.
+Formally, let denote a natural language message, and denote a cover object.
+The encoder embeds into , producing the stego object .
+The corresponding decoder aims to recover the original message as .
+The semantic steganography task must satisfy the following constraints:
+Specifically, recoverability measures the semantic consistency of the recovery message using the similarity function .
+Imperceptibility evaluates the similarity between cover and stego objects, where measures perceptual distortion, and denotes the maximum allowable perceptual difference.
+Security reflects the indistinguishability between cover and stego objects against steganalysis.
+represents the output of the steganalysis tool.
+In essence, semantic steganography generalizes the notion of information hiding from the signal level to the semantic level.
+By encoding linguistic meaning rather than raw bits, it establishes a unified framework for embedding natural language semantics across arbitrary modalities.
 3.2 Benchmark
@@
 We divide the messages into three levels of granularity: short messages (one sentence), medium messages (two to three sentences), and long messages (full articles), denoted as IVT-S, IVT-M, and IVT-L, respectively.
-Please refer to Appendix A for more details about the IVT dataset.
-4 Our Method
-4.1 Representation Information as Tokens
-The representation of secret text is a critical aspect of semantic steganography.
-Previous methods convert each character of a word into bits using UTF encoding[tancik2020stegastamp, zhu2018hidden, wengrowski2019light].
-However, these primitive approaches have two significant limitations.
-First, UTF encoding lacks semantic information, which is essential for representing sentence-level content.
-Second, representing a sentence as a bitstream is inefficient, as the resulting sequence becomes excessively long and difficult to process.
-We hypothesize that such primitive representations are particularly fragile when encoding richer semantic content.
+Please refer to Sec.Β 7 of supplementary materials for more details.
+3.3 Sentence-to-Image Steganography
+Specifically, we formalize the sentence-to-image steganography task as a concrete instantiation of the semantic steganography.
+Following the definition above, the object in this task is specified as an image .
+The objective is to secretly embed natural-language messages within the image while ensuring that the hidden information can be accurately recovered.
+4 Method
+4.1 Representing Information as Tokens
+The representation of secret message is a critical aspect of semantic steganography.
+Previous methods convert each character of a word into bits using UTF encoding[60, 83, 68].
+However, such character-level representations have two significant limitations.
+First, UTF encoding lacks semantic compression capability, treating each character equally regardless of its meaning.
+Second, encoding a sentence as a bitstream leads to an excessively long sequence that is difficult to process.
+We hypothesize that such primitive representations are fragile when encoding richer semantic content.
 Therefore, it is necessary to explore more effective representations for sentence-level information.
-To address these issues, we leverage the tokenizer of the LLM to convert the secret sentence into tokens.
-Tokenization is a widely used technique in NLP, where sentences are broken down into smaller units called Tokens.
-Tokens capture richer semantic information and facilitate modeling relationships between words.
-Moreover, tokenization significantly reduces sequence length, which benefits downstream processing.
-Using the tokenizer, we represent the secret sentence as a sequence of tokens.
-This approach not only preserves the semantic structure of the sentence but also ensures an efficient representation for embedding.
+To address these issues, we leverage the tokenizer to convert the secret sentence into tokens.
+Tokenization [67] is a widely used technique in NLP, where sentences are broken down into smaller units called Tokens.
+Unlike character-level encoding, tokens correspond to semantically meaningful units, such as words, sub-words, or phrases, allowing a more compact representation.
+This not only preserves the semantic structure of the sentence but also significantly shortens the sequence length, improving the efficiency and stability of downstream embedding and decoding processes.
 4.2 LLMs for Steganography
-As shown in FigureΒ 2 the consists of three key modules: a pre-trained LLM for embedding and decoding the secret information, a Token-to-Patch MLP for aligning text features and image feature, and a Patch-to-Token MLP for embedding visual features to token embeddings.
+As shown in Fig.Β 2, the consists of three key modules: a pre-trained LLM for embedding and decoding the secret information, a Token-to-Patch MLP for aligning text features and image features, and a Patch-to-Token MLP for embedding visual features to token embeddings.
 Notably, is a universal framework in which the LLM can be ANY pre-trained LLM.
@@
 Recent studies, however, have demonstrated that LLM can serve as a universal interface with different instructions in language to complete different tasks.
-In the computer vision community, researchers have explored integrating LLMs with vision backbones for segmentation, localization, and other visual tasks.
+In the computer vision community, researchers have explored integrating LLMs with vision backbones for segmentation [35], grounding [69, 50, 30], and other visual tasks [74].
 Inspired by these works, we employ an LLM to generate embeddings conditioned by secret information, which will be inserted into the cover image to produce the stego image.
 Owing to the strong generalization ability of LLMs, we are able to compress the semantic information into a few embeddings.
-Specifically, as illustrated in Figure 2, we design an embedding prompt and concatenate it with a secret message to instruct the LLM to embed the secret message into embedding, which we call secret message embeddings (SMEs).
+Specifically, as illustrated in Fig.Β 2, we design an embedding prompt and concatenate it with a secret message to instruct the LLM to embed the secret message into embeddings, which we call secret message embeddings (SMEs).
 We call this strategy Prompt-Guide Embedding.
@@
 To insert SMEs into the cover image, we use to align the token embeddings with the image features.
-Then, we adapt the UDH method [zhang2020udh], which directly adds the SMEs to the cover image to produce a stego image.
+Then, we adapt the UDH method [75], which directly adds the SMEs to the cover image to produce a stego image.
 Previous methods generate the stego image conditioned on both the cover image and the secret message.
@@
 Finally, we get the recovery message from the LLM.
-Template designing As illustrated in Figure 3, we carefully design the embedding and decoding prompts to instruct the LLM to embed and recover the secret message.
+Template Designing.
+As illustrated in Fig.Β 3, we carefully design the embedding and decoding prompts to instruct the LLM to embed and recover the secret message.
 We introduce two special tokens: and , which are used to separate secret messages from the prompt.
@@
 Trainable Parameters.
-To preserve the knowledge of the LLM, we use LoRA [hu2022lora] for fine-tuning.
+To preserve the knowledge of the LLM, we use LoRA [26] for fine-tuning.
 Additionally, because we introduce new tokens into the vocabulary (e.g.
-), we also set the LLM token embeddings (), the LLM head (), and the projection layer trainable.
+), we also set the LLM token embeddings () and the LLM head () trainable.
 By the way, the and are also trainable.
+| Methods | IVT-S | IVT-M | IVT-L | |||||||||||||||
+| Secret/Recovery | Cover/Stego | Secret/Recovery | Cover/Stego | Secret/Recovery | Cover/Stego | |||||||||||||
+| WER | BLEU | ROUGE | BERT-S | PSNR | SSIM | WER | BLEU | ROUGE | BERT-S | PSNR | SSIM | WER | BLEU | ROUGE | BERT-S | PSNR | SSIM | |
+| StegaStamp [60] | 0.452 | 0.241 | 0.138 | 0.347 | 35.6 | 0.927 | - | - | - | - | - | - | - | - | - | - | - | - |
+[…]
+| FPGP [78] | 0.150 | 0.882 | 0.925 | 0.914 | 43.5 | 0.954 | 0.758 | 0.462 | 0.539 | 0.538 | 35.6 | 0.874 | - | - | - | - | - | - |
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
 Stage 1: Pre-Training for Steganography.
-As shown in FigureΒ 4, the training process consists of two stages.
-In the first stage, we train the LLM with LoRA, Token-to-Patch MLP , and Patch-to-Token MLP .
+As shown in Fig.Β 4, the training process consists of two stages.
+In the first stage, we train the LLM (with LoRA), together with and .
 In this stage, we DO NOT introduce the cover image.
-We wrap the secret message with and to construct the input sequence .
+We wrap the secret message with and to construct the input sequence :
 Given the embedding prompt along with , we feed them into the LLM, which outputs a response token sequence , which can be formulated as:
@@
 Finally, we concatenate it with decode prompt and feed it into the LLM to obtain the recovery message , which can be formulated as:
-Stage 2: fine-tuning in stego image.
-In the second stage, we FREEZE the LLM and the , TRAIN the , and introduce the cover image .
+Stage 2: Fine-Tuning on Stego Images.
+In the second stage, we FREEZE the LLM and , TRAIN , and introduce the cover image .
 After obtaining , we reshape it from patch representation to the image representation , and add it to to obtain stego image .
-The entire process can be formulated as follow:
+The entire process can be formulated as follows:
 After receiving the stego image, we divide the it into patches to transform it into a 1D sequence and feed it into to obtain .
 Finally, we concatenate decode prompt with the and feed it into the LLM to get the recovery message , which can be formulated as follows:
-Training objects. The is trained using the text generation loss and the embedding loss . The overall objective of each stage is defined as the weighted sum of these losses:
-Specifically, is the auto-regressive cross-entropy (CE) loss for text generation, and is the L1 loss, which encourages the model to introduce minimal artifacts in the cover image.
-Those losses can be formulated as:
+Training Objectives. The is trained using the text generation loss and the embedding loss . The overall objective of each stage is defined as the weighted sum of these losses:
+Specifically, is the cross-entropy (CE) loss for text generation, and is the L1 loss, which encourages the model to introduce minimal artifacts in the cover image. Those losses can be formulated as:
 5 Experiments
@@
 is a universal framework compatible with any pre-trained LLM.
-In this paper, we instantiate it with four models: Llama3.2-1B [grattafiori2024llama], Qwen2.5-0.5B [qwen2025qwen25technicalreport], Gemma3-1B [team2025gemma], and MiniCPM-1B [hu2024minicpm].
-Implementation details.
-We employ AdamW optimizer [loshchilov2017decoupled].
-The weights of the text generation loss and the embedding loss are set to 1.0 and 1.0, respectively.
-More details can be found on the Appendix B.1.
+In this paper, we instantiate it with four models: Llama3.2-1B [21], Qwen2.5-0.5B [55], Gemma3-1B [61], and MiniCPM-1B [28].
+We chose edge-deployable language models with parameters fewer than 1B, which guarantees lightweight operation while maintaining the scalability to larger models.
+Implementation Details. We employ AdamW optimizer [44]. The weights and are set to 1.0 and 1.0, respectively. More details can be found in Sec.Β 8.2 of the supplementary material.
 Datasets.
-For the secret message, we construct the training dataset using the CNN News Stories dataset [cnn_stories_dataset], news headlines [see2017get, ao-etal-2021-pens], and paper titles [farhangi2022protoformer].
-For the cover image, we use the COCO training dataset [lin2014microsoft].
-Please refer to Appendix B.2 for detailed information about training datasets.
-We evaluate the performance of steganographic models on the sentence-to-image steganography task using the IVT dataset.
-In experiments, we use the COCO test dataset as the cover image dataset.
-For more details about IVT, please refer to Appendix A.
+For the secret message, we use Wanjuan 1.0 [25], a comprehensive corpus designed for large language model pre-training.
+We randomly select 320,000 samples with fewer than 1,000 words as secret messages for training.
+For the cover image, we use the COCO training dataset [41].
+Please refer to the supplementary material for detailed information about the training dataset.
+We evaluate the performance of steganographic models on the sentence-to-image steganography task, using the IVT dataset as secret messages and the COCO test dataset as cover images.
+For more details about IVT, please refer to Sec.Β 7 in the supplementary material.
 Evaluation Metrics.
-To measure the quality of the stego image, we use PSNR and SSIM to evaluate embedding performance.
-To assess recovery performance, we use word error rate (WER), BLEU-4 [papineni2002bleu], and ROUGE-L [lin2004rouge], and BERT-Score [zhang2019bertscore] to evaluate the semantic similarity between the ground truth and the recovery message.
-Benchmarks. We compare with different methods, including DwtDct [Chen2009DigitalWA], StegaStamp [tancik2020stegastamp] and LanNet [lan2023robust]. We train StegaStamp and LanNet from scratch.
+The performance of the steganographic model is evaluated from three complementary perspectives: recoverability, imperceptibility, and security.
+For recoverability, we evaluate the recovery text using Word Error Rate (WER), BLEU-4 [47], ROUGE-L [40], and BERT-Score [77], which jointly assess lexical accuracy and semantic consistency.
+For imperceptibility, we adopt PSNR and SSIM to measure the visual similarity between the cover and stego images.
+For security, we employ state-of-the-art steganalysis tools including SRNet [8], SiaSigNet [71], and DBS2Net [27], to quantify the detectability of stego images.
 5.2 Sentence-to-Image Steganography Results
-Sentence-to-Image results are shown in Table 1. Notably, while existing methods struggle with this task, our demonstrates strong performance.
+Sentence-to-Image results are shown in Tab.Β 1. Notably, while existing methods struggle with this task, our demonstrates strong performance.
 Results of Existing Methods.
-Notably, existing methods are not suitable for the sentence-to-image steganography task, which fundamentally differs from watermarking in that it requires encoding a substantial amount of semantic information.
-As shown in TableΒ 1, StegaStamp (512-bit capacity) fails to recover secret messages in IVT-S, while DwtDct (768-bit capacity) performs well only on IVT-S but fails to recover messages in IVT-M, which exceed its capacity, not to mention handling IVT-L.
-However, LanNet (1024-bit capacity) also fails on IVT-L.
+Existing methods are not suitable for the sentence-to-image steganography task, which fundamentally differs in that it requires encoding a substantial amount of semantic information.
+As shown in Tab.Β 1, StegaStamp fails to recover secret messages in IVT-S, while DwtDct and FPGP perform well only on IVT-S, but fail to recover messages in IVT-M, not to mention handling IVT-L.
+However, LanNet and CRMark also fail on IVT-L.
+To ensure sufficient capacity for the IVT dataset, we retrain all of the baselines and increase their steganographic capacity for each IVT dataset.
+However, we observe a severe mode collapse problem, which is consistent with the phenomenon reported in [52].
+We provide a detailed analysis in Sec.Β 12.2 of the supplementary material.
+This issue primarily stems from limitations in the architectural design and bitstream representation of secret information.
+As a result, these models are restricted to hiding bit-level information and cannot handle semantic-level content, such as sentences.
 Results of . Existing methods fail to address this task effectively, whereas successfully handles secret messages containing complex and semantically rich natural language information.
-On the IVT-S, all four variants demonstrate competitive performance compared to traditional baselines in secret recovery while significantly outperforming them in visual quality of stego image.
+On the IVT-S, all four variants demonstrate competitive performance compared to traditional baselines in secret recovery while significantly outperforming them in the visual quality of the stego image.
 Specifically, -MiniCPM-1B achieves the best overall performance, while -Qwen2.5-0.5B, -Llama3.2-1B, and -Gemma3-1B also deliver competitive results.
-On the IVT-M, which involves longer and more complex message, -MiniCPM-1B maintains superior performance across most evaluation metrics.
--Qwen2.5-0.5B and -Llama3.2-1B show slight performance degradation, while -Gemma3-1B demonstrates inferior semantic accuracy despite achieving high image quality.
+| Capacity (Token) | Compress Ratio | Secret / Recovery | Cover / Stego | ||||
+| WER | BLEU-4 | ROUGE-L | BERT-Score | PSNR | SSIM | ||
+| 32 | 1:2 | 0.033 | 0.947 | 0.975 | 0.971 | 34.6 | 0.875 |
+| 64 | 1:1 | 0.035 | 0.952 | 0.977 | 0.969 | 31.6 | 0.773 |
+| 128 | 2:1 | 0.085 | 0.892 | 0.947 | 0.934 | 33.5 | 0.853 |
+| 256 | 4:1 | 0.214 | 0.725 | 0.860 | 0.858 | 32.4 | 0.809 |
+| 512 | 8:1 | 0.875 | 0.237 | 0.474 | 0.708 | 31.5 | 0.770 |
+On the IVT-M, which involves longer and more complex messages, our variants significantly outperform prior methods.
+-Qwen2.5-0.5B achieves the best recovery performance among all variants while maintaining high visual quality.
+-MiniCPM-1B and -Gemma3-1B exhibit comparable performance.
 On the IVT-L, all models exhibit performance degradation compared to their results on IVT-M.
--Qwen2.5-0.5B get the top performance, -MiniCPM-1B and -Llama3.2-1B demonstrate moderate performance declines, and -Gemma3-1B exhibits the lowest semantic accuracy.
+-Qwen2.5-0.5B get the top recovery performance, -MiniCPM-1B and -Llama3.2-1B demonstrate moderate performance declines, and -Gemma3-1B exhibits the lowest semantic accuracy.
 Experimental results demonstrate that all variants achieve strong performance on both IVT-S and IVT-M.
 However, for these models, IVT-L remains challenging, which comprises multi-sentence structures characterized by complex syntactic patterns and diverse lexical content.
-More details can be found in Appendix E.
-| IVT-S | IVT-M | IVT-L | ||||||||||||||||
-| Methods | Secret/Recovery | Cover/Stego | Secret/Recovery | Cover/Stego | Secret/Recovery | Cover/Stego | ||||||||||||
-| WER | BLEU | ROUGE | BERT-S | PSNR | SSIM | WER | BLEU | ROUGE | BERT-S | PSNR | SSIM | WER | BLEU | ROUGE | BERT-S | PSNR | SSIM | |
-| StegaStamp [tancik2020stegastamp] | 0.452 | 0.241 | 0.138 | 0.347 | 35.6 | 0.927 | - | - | - | - | - | - | - | - | - | - | - | - |
-| DwtDct [chen2006dwt] | 0.007 | 0.699 | 0.996 | 0.996 | 31.8 | 0.898 | 0.701 | 0.490 | 0.713 | 0.829 | 27.7 | 0.776 | - | - | - | - | - | - |
-| LanNet [lan2023robust] | 0.004 | 0.706 | 0.997 | 0.996 | 42.1 | 0.951 | 0.244 | 0.698 | 0.836 | 0.857 | 41.3 | 0.949 | 0.824 | 0.031 | 0.288 | 0.630 | 40.9 | 0.941 |
-| S2LM-Qwen2.5-0.5B | 0.045 | 0.619 | 0.873 | 0.919 | 52.1 | 0.997 | 0.050 | 0.906 | 0.944 | 0.952 | 48.0 | 0.992 | 0.112 | 0.840 | 0.901 | 0.913 | 39.0 | 0.951 |
-| S2LM-MiniCPM-1B | 0.046 | 0.635 | 0.923 | 0.959 | 59.8 | 0.999 | 0.061 | 0.913 | 0.962 | 0.965 | 52.6 | 0.996 | 0.197 | 0.762 | 0.859 | 0.889 | 46.9 | 0.988 |
-| S2LM-Llama3.2-1B | 0.049 | 0.616 | 0.874 | 0.927 | 53.2 | 0.998 | 0.080 | 0.840 | 0.886 | 0.908 | 48.0 | 0.993 | 0.171 | 0.653 | 0.743 | 0.798 | 40.8 | 0.973 |
-| S2LM-Gemma3-1B | 0.100 | 0.498 | 0.746 | 0.834 | 53.2 | 0.998 | 0.141 | 0.745 | 0.832 | 0.871 | 51.7 | 0.997 | 0.254 | 0.678 | 0.795 | 0.840 | 44.8 | 0.984 |
+Additional robustness experiments demonstrate that exhibits strong resistance to various attacks while maintaining a remarkably high capacity.
+More details can be found in Sec.Β 12.4 of the supplementary material.
 Qualitative Results.
-As depicted in Figure 5, we provide a visualization of stego images and recovery results of -Qwen2.5-0.5B in different length of information.
+As depicted in Fig.Β 5, we provide a visualization of stego images and recovery results of -Qwen2.5-0.5B in different lengths of information.
 We mark the difference in red.
-As we can see, the recovery message and the ground truth keep high semantic similarity, while the stego image still has a good visual quality.
-More samples can be found in Appendix I.
-| Capacity | Compress Ratio | Secret / Recovery | Cover / Stego | ||||
-| WER | BLEU-4 | ROUGE-L | BERT-Score | PSNR | SSIM | ||
-| 32 Token | 1:2 | 0.0331 | 0.9470 | 0.9752 | 0.9711 | 34.61 | 0.8759 |
-| 64 Token | 1:1 | 0.0353 | 0.9523 | 0.9779 | 0.9694 | 31.64 | 0.7735 |
-| 128 Token | 2:1 | 0.0855 | 0.8920 | 0.9476 | 0.9347 | 33.57 | 0.8536 |
-| 256 Token | 4:1 | 0.4508 | 0.4847 | 0.7074 | 0.7454 | 29.53 | 0.5802 |
-5.3 Capacity of the Cover Image
-In this section, we conduct empirical experiments to investigate the capacity of the cover image under the framework.
-We analyze the capacity in terms of tokens, setting the secret message length to 32, 64, 128, and 256 tokens, while fixing the cover image resolution at .
-We train -Qwen2.5-0.5B from scratch for each setting.
-The cover image is divided into 64 patches, which is equal to the number of secret message embeddings (SMEs).
-For clearer analysis, we define the compression ratio as the number of SMEs divided by the number of patches.
-| Two-Stage Training | Prompt | Embedding Loss | Mask Strategy | Secret Text / Recovery Text | Cover / Stego | ||||
-| WER | BLEU-4 | ROUGE-L | BERT-Score | PSNR | SSIM | ||||
-| 0.3162 | 0.5925 | 0.6781 | 0.6816 | 34.13 | 0.8156 | ||||
-| β | 0.2210 | 0.6819 | 0.7403 | 0.8051 | 32.93 | 0.8460 | |||
+As we can see, the recovery message and the ground truth keep a high semantic similarity, while the stego image still has a good visual quality.
+More samples can be found in Sec.Β 15 of the supplementary material.
+| No. | Two-Stage Training | Prompt | Embedding Loss | Mask Strategy | Secret / Recovery | Cover / Stego | ||||
+| WER | BLEU-4 | ROUGE-L | BERT-S | PSNR | SSIM | |||||
+| 1 | 0.3162 | 0.5925 | 0.6781 | 0.6816 | 34.13 | 0.8156 | ||||
+| 2 | β | 0.2210 | 0.6819 | 0.7403 | 0.8051 | 32.93 | 0.8460 | |||
 […]
 […]
-| β | β | β | 0.2379 | 0.6703 | 0.7496 | 0.8119 | 34.24 | 0.7820 | |
+| 5 | β | β | β | 0.2379 | 0.6703 | 0.7496 | 0.8119 | 34.24 | 0.7820 | |
 […]
-As illustrated in TableΒ 2, the model maintains strong decoding performance at 32 and 64, 128 tokens, with BERT-Scores above 0.9.
-However, at 256 tokens, frequent decoding errors emerging.
-These results indicate that while the model has effective information compression ability, its capacity saturates beyond a certain threshold, leading to a noticeable degradation in recovery quality.
-Based on these empirical findings, we propose a practical heuristic: each image patch can stably carry approximately two tokens without significant loss in decoding fidelity.
-This guideline provides an interpretable measure of capacity for future designs and scaling strategies, offering a concrete reference point for balancing payload length and image quality in sentence-to-image steganography.
+5.3 Capacity Analysis
+In this section, we conduct empirical experiments to investigate the capacity of the cover image under the framework.
+We analyze the capacity in terms of tokens, setting the secret message length to 32, 64, 128, 256, and 512 tokens, while fixing the cover image resolution at .
+We train -Qwen2.5-0.5B from scratch for each setting.
+The cover image is divided into 64 patches, which is equal to the number of SMEs.
+For more precise analysis, we define the compression ratio as the number of message tokens divided by the number of patches.
+As illustrated in Tab.Β 2, the model maintains strong decoding performance for message lengths ranging from 32 to 256 tokens, with BERT-Scores above 0.8.
+However, at 512 tokens, frequent decoding errors emerge.
+These results indicate that while the has the information compression ability, its capacity saturates beyond a certain threshold, leading to a noticeable degradation in recovery quality.
+Based on these empirical findings, we propose a practical heuristic: each image patch can stably carry approximately four tokens without significant loss in decoding fidelity.
+This guideline provides an interpretable measure of capacity for future designs and scaling strategies, offering a concrete reference point for balancing payload and image quality in sentence-to-image steganography.
+5.4 Steganalysis
+Following previous works [23, 29, 31, 83, 75], we gradually increase the number of training samples to investigate the minimum number of leaking samples required to train steganalysis models for detecting stego images.
+We retain the SRNet, SiaStegNet, and DBS2Net with cover/stego image pairs on the COCO dataset.
+As shown in Fig.Β 6, with the increasing number of training samples, the still achieves lower detection accuracy.
+More details about the statistical steganalysis can be found in Sec.Β 9 of the supplementary material.
 5.5 Ablation Study
@@
 Effectiveness of Two-Stage Training.
-As shown in Table 3, without two-stage training strategy, the struggles to recover the secret message.
-That may be because introducing the image modality at an early stage disrupts the LLMβs ability to learn steganographic mechanisms.
+Without the two-stage training strategy, the struggles to recover the secret message (No.1 vs.
+No.2, No.5 vs.
+No.6).
+That may be because introducing the image modality at an early stage disrupts the LLM to learn steganographic mechanisms.
 Effectiveness of the Prompt.
@@
 Similarly, during decoding, we directly feed the features extracted by the Patch-to-Token MLP into the LLM to recover the original message.
-As shown in TableΒ 3, this approach leads to performance improvements of 7.3%, 7.5%, and 3.0% in BLEU-4, ROUGE-L, and BERT-Score, respectively.
+Comparing experiments No.2 and No.3 reveals that this approach leads to performance improvements of 7.3%, 7.5%, and 3.0% in BLEU-4, ROUGE-L, and BERT-Score, respectively.
 These results confirm that the prompt effectively guides the LLM in understanding and executing the sentence-to-image task.
 Effectiveness of Embedding Loss.
-In the TableΒ 3, we show the contribution of the embedding loss .
-The is used to encourage the LLM to produce minimal artifacts in the cover image and is applied only in the first stage.
-However, experiment result shows that the introduction of the can improve the embedding quality with improvement of 15.61 dB and 0.165 in the PSNR and SSIM.
+A comparison between experiments No.3 and No.4 reveals the contribution of .
+is used to encourage the LLM to produce minimal artifacts in the cover image and is applied only in the first stage.
+The experimental results show that the introduction of can improve the embedding quality, with gains of 15.61 dB and 0.165 in the PSNR and SSIM.
 Effectiveness of Mask Strategy.
-In decoding process of the first stage, we mask some input tokens to encourage LLM distribute the information in the whole cover image.
-As shown in FigureΒ 6, the SMEs are evenly distributed after training with the mask strategy, and significant improvements in recovery quality are observed in TableΒ 3.
+In the decoding process of the first stage, we mask some input tokens to encourage the LLM to distribute information across the entire cover image.
+As shown by the comparison between experiments No.4 and No.6, this strategy leads to significant improvements in recovery quality.
+We visualize the SMEs in Sec.Β 11.1 of the supplementary material, which shows that the SMEs are evenly distributed with the mask strategy.
 6 Conclusion
-In this work, we propose a new steganographic task, sentence-to-image steganography, which is an instance of semantic steganography.
-Additionally, we present an evaluation benchmark, Invisible Text (IVT), comprising a diverse set of sentence-level texts drawn from existing datasets and generated by large language models (LLMs).
-Finally, we present , a novel steganographic framework that leverages LLMs to embed sentence-level information into images.
-We hope our work sheds new light on the advancement of semantic steganography.
-Appendix A Invisible Text Dataset
+In this work, we propose a novel concept of semantic steganography, which aims to hide high-level semantic information in carriers and redefines the traditional paradigm of steganography.
+Building on this insight, we present sentence-to-image steganography as an instance, which focuses on embedding sentence-level information into images.
+To achieve this task, we propose , a novel steganographic framework that leverages LLMs to embed sentence-level information into images, achieving a feat that previously has been out of reach.
+Additionally, we present an evaluation benchmark, Invisible Text (IVT), comprising a diverse set of sentence-level texts drawn from existing datasets and generated by LLMs.
+We hope our work sheds new light on the advancement of semantic steganography for the entire community.
+Supplementary Material
+Overview of the Supplementary Material
+This supplementary document provides a comprehensive extension to the main paper, organized as follows:
+β’
+Dataset Details (Sec.Β 7): We introduce the proposed Invisible Text (IVT) benchmark, describe the construction of its three granular subsets (IVT-S/M/L), and provide detailed statistics on their text length, bit length, domains, and licenses.
+We also describe the construction of the LLM-generated IVTG dataset.
+β’
+Experimental Details (Sec.Β 8): We provide full training configurations for and baselines, including training dataset construction, model variants, optimization settings, implementation details, and retraining protocols for comparison methods.
+β’
+Statistical Steganalysis (Sec.Β 9): We conduct statistical steganalysis using StegExpose to evaluate the detectability of stego images and show that our method remains highly secure against classical statistical attacks.
+β’
+Additional Experimental Results (Sec.Β 10): We present comprehensive supplementary experiments, including using vision-language models within , constructing and benchmarking on the LLM-generated IVTG dataset, testing non-semantic messages, analyzing decoding with the original and cross-instance LLMs, applying 8-bit quantization on , and providing additional qualitative results.
+β’
+Ablation Study Details (Sec.Β 11): We visualize the spatial distribution of secret message embeddings under the mask strategy to analyze how it shapes SME layouts.
+β’
+Analysis of Existing Methods (Sec.Β 12): We analyze how existing UTF-based bitstream representations and MLP designs inherently constrain capacity and scalability of current steganographic methods.
+β’
+Discussion (Sec.Β 13): We discuss the motivation behind semantic steganography, as well as the broader impacts, potential risks, and mitigation strategies associated with .
+β’
+Background and Preliminaries (Sec.Β 14): We clarify the terminology and related notions, such as the cover image and the stego image, used throughout this work.
+7 Invisible Text Dataset
 To evaluate the effectiveness of semantic steganography, we construct a new benchmark dataset named InVisible Text (IVT).
@@
 This three-granularity design enables comprehensive performance analysis across varying semantic loads, offering insights into the scalability, robustness, and text fidelity of semantic steganographic frameworks.
-A.1 IVT-S
+| Category | Source | Bit Length | Word Length | Unique Word | Sample Number | License | ||||
+| Average | Max | Min | Average | Max | Min | |||||
+| Paper title | arxiv-10 [16] | 592.2 | 1296 | 120 | 9.7 | 20 | 2 | 4576 | 1000 | GPL-3.0 |
+| News title | PENS [3] | 506.9 | 1072 | 120 | 10.3 | 20 | 2 | 5506 | 1000 | MSR License |
+| Article title | examiner [34] | 423.1 | 800 | 40 | 8.6 | 19 | 1 | 4873 | 1000 | CC0 1.0 License |
+| People name | name&code [63] | 113.8 | 200 | 72 | 2.1 | 4 | 2 | 857 | 1000 | - |
+| Sentence | cmn-en [32] | 247.3 | 848 | 40 | 6.2 | 18 | 1 | 1946 | 1000 | CC-BY 2.0 License |
+| Dialogue | SP-S1 [1] | 202.9 | 520 | 32 | 4.5 | 9 | 1 | 1924 | 1000 | - |
+| Comment | sentiment140 [20] | 486.2 | 1120 | 56 | 10.7 | 20 | 1 | 4374 | 1000 | - |
+| Overall | - | 367.5 | 836.6 | 68.6 | 7.4 | 15.7 | 1.4 | 19941 | 7000 | - |
+| Category | Source | Bit Length | Word Length | Unique Word | Sample Number | License | ||||
+| Average | Max | Min | Average | Max | Min | |||||
+| Paper title | arxiv-10 [16] | 1210.0 | 1904 | 696 | 22.0 | 38 | 20 | 7776 | 1000 | GPL-3.0 |
+| New title | PENS [3] | 1369.4 | 1968 | 624 | 23.0 | 41 | 20 | 8003 | 1000 | MSR License |
+| News highlight | CNN and Daily mail [56] | 2303.6 | 4896 | 736 | 50.3 | 100 | 20 | 13119 | 1000 | Apache 2.0 |
+| Translate text | WMT-14 [7] | 1610.1 | 4776 | 672 | 33.8 | 98 | 20 | 6760 | 1000 | - |
+| Joke story | stupid joke [53] | 2042.1 | 17648 | 592 | 47.6 | 100 | 20 | 10637 | 1000 | - |
+| Translate text | europarl [33] | 1618.8 | 5344 | 744 | 34.0 | 99 | 20 | 6735 | 1000 | - |
+| Dialogue | South Park [1] | 1367.1 | 4040 | 688 | 31.7 | 92 | 20 | 3896 | 1000 | - |
+| Comment | hate text [13] | 989.9 | 1120 | 552 | 23.4 | 33 | 20 | 6564 | 1000 | MIT License |
+| Moral | moral story [22] | 1369.4 | 4728 | 680 | 32.5 | 97 | 20 | 3179 | 1000 | CC BY-NC-SA 3.0 |
+| Overall | - | 1542.3 | 17648 | 552 | 33.1 | 100 | 20 | 49423 | 9000 | - |
+| Category | Source | Bit Length | Word Length | Unique Word | Sample Number | License | ||||
+| Average | Max | Min | Average | Max | Min | |||||
+| Paper abstract | arxiv-10 [16] | 8589.8 | 15360 | 2424 | 157.8 | 325 | 50 | 22705 | 1000 | GPL-3.0 |
+| Wiki abstract | wiki [18] | 7840.5 | 40776 | 2216 | 157.5 | 863 | 50 | 26582 | 1000 | GFDL/CC BY-SA |
+| Story | stupid joke [53] | 7285.7 | 41168 | 1920 | 166.4 | 997 | 50 | 25232 | 1000 | - |
+| Story | moral story [22] | 11374.1 | 42576 | 1920 | 265.0 | 956 | 50 | 20898 | 1000 | - |
+| Story | Tiny Story [15] | 6957.5 | 33192 | 2600 | 166.4 | 813 | 64 | 10348 | 1000 | - |
+| User review | IMDB [49] | 10457.4 | 45736 | 2048 | 231.0 | 1000 | 50 | 34259 | 1000 | IMDB License |
+| User review | yelp [4] | 5996.6 | 39664 | 1968 | 138.9 | 871 | 50 | 18029 | 1000 | Yelp License |
+| Composition | asap-aes [17] | 10435.9 | 39216 | 1912 | 238.5 | 897 | 50 | 17289 | 1000 | - |
+| Overall | - | 8675.3 | 45736 | 1912 | 194.5 | 1000 | 50 | 118472 | 8000 | - |
+7.1 IVT-S
 IVT-S is the simplest variant in IVT, focusing on short, low-capacity messages.
@@
 IVT-S serves as a lightweight benchmark for evaluating models under low-information embedding scenarios.
-The IVT-S dataset covers a wide range of short textual types, including real person names [name_country_origin_dataset], news headlines [ao-etal-2021-pens], paper titles [farhangi2022protoformer], Twitter comments [go2009twitter], and dialogue snippets [southparkdata].
-For more details on IVT-S, please refer to Table 4.
-| Category | Source |
-License | ||||||||
-| paper title | arxiv-10 [farhangi2022protoformer] | 9.7 | 592.2 | 4576 | 1000 | GPL-3.0 | ||||||||
-| news title | PENS [ao-etal-2021-pens] | 10.3 | 506.9 | 5506 | 1000 | MSR License | ||||||||
-| article title | examiner [examine_examiner_dataset] | 8.6 | 423.1 | 4873 | 1000 | CC0 1.0 | ||||||||
-| people name | name&code [name_country_origin_dataset] | 2.1 | 113.8 | 857 | 1000 | - | ||||||||
-| sentence | cmn_en [manythings_anki_dataset] | 6.2 | 247.3 | 1946 | 1000 | CC-BY 2.0 | ||||||||
-| dialogue | South Park [southparkdata] | 4.5 | 202.9 | 1924 | 1000 | - | ||||||||
-| comment | sentiment140 [go2009twitter] | 10.7 | 486.2 | 4374 | 1000 | - | ||||||||
-| overall | - | 7.4 | 367.5 | 19941 | 7000 | - |
-A.2 IVT-M
+7.2 IVT-M
 IVT-M is a medium-granularity subset in IVT, positioned between the short messages of IVT-S and the full-paragraph texts of IVT-L.
@@
 IVT-M reflects more realistic use cases than IVT-S by requiring the model to encode and recover messages with non-trivial structure and meaning.
-Messages in IVT-M are sourced from a variety of domains, including CNN news abstracts [see-etal-2017-get], Wikipedia summaries [wikidump], IMDB reviews [imdb_reviews_dataset].
-For more details about IVT-M, please refer to Table 5.
-| Category | Source |
-License | ||||||||
-| paper title | arxiv-10 [farhangi2022protoformer] | 22.0 | 1210.0 | 7776 | 1000 | GPL-3.0 | ||||||||
-| new title | PENS [ao-etal-2021-pens] | 23.0 | 1369.4 | 8003 | 1000 | MSR License | ||||||||
-| news highlight | CNN and Daily mail [see2017get] | 50.3 | 2303.6 | 13119 | 1000 | Apache 2.0 | ||||||||
-| translate text | WMT-14 [bojar-EtAl:2014:W14-33] | 33.8 | 1610.1 | 6760 | 1000 | - | ||||||||
-| joke story | stupid joke [pungas] | 47.6 | 2042.1 | 10637 | 1000 | - | ||||||||
-| translate text | europarl [koehn2005europarl] | 34.0 | 1618.8 | 6735 | 1000 | - | ||||||||
-| dialogue | South Park [southparkdata] | 31.7 | 1367.1 | 3896 | 1000 | - | ||||||||
-| comment | hate text [hateoffensive] | 23.4 | 989.9 | 6564 | 1000 | MIT | ||||||||
-| moral | moral story [guan2022corpus] | 32.5 | 1369.4 | 3179 | 1000 | CC BY-NC-SA 3.0 | ||||||||
-| overall | - | 33.1 | 1542.3 | 42923 | 9000 | - |
-A.3 IVT-L
+7.3 IVT-L
 IVT-L represents the largest and most challenging subset of the IVT dataset, designed to evaluate a modelβs ability to embed and recover long-form semantic content.
-Each secret message in IVT-L consists of a full paragraph containing multiple sentences with coherent structure, complex syntax, and a diverse vocabulary.
+Each secret message in IVT-L consists of a full paragraph containing multiple sentences with coherent structure, complex syntax, and a diverse vocabulary, with a length of about 100 to 500 words.
 Compared to IVT-S and IVT-M, IVT-L significantly enhances the semantic richness and information density of the hidden text, thereby imposing greater demands on a modelβs capacity, fidelity, and ability to preserve semantics.
 This subset serves as a benchmark for evaluating whether steganographic models can scale to real-world scenarios that involve concealing full-length, meaningful textual content.
-IVT-L includes long-form messages drawn from a variety of sources, including narrative news passages, short stories [guan2022corpus, eldan2023tinystories], research paper abstracts [farhangi2022protoformer], IMDB reviews [imdb_reviews_dataset].
+IVT-L includes long-form messages drawn from a variety of sources, including narrative news passages, short stories [22, 15], research paper abstracts [16], IMDB reviews [49].
 These samples span multiple domains and writing styles, providing a comprehensive benchmark for evaluating sentence-to-image steganography in high-capacity settings.
 For more details about IVT-L, please refer to Table 6.
-| Category | Source |
-License | ||||||||
-| paper abstract | arxiv-10 [farhangi2022protoformer] | 156.55 | 8512.75 | 22550 | 1000 | GPL-3.0 | ||||||||
-| wiki abstract | wiki [wikidump] | 140.67 | 6962.01 | 25117 | 1000 | GFDL/CC BY-SA | ||||||||
-| story | stupid joke [pungas] | 153.37 | 6725.4 | 23701 | 1000 | - | ||||||||
-| story | moral story [guan2022corpus] | 225.92 | 9679.95 | 18412 | 1000 | - | ||||||||
-| story | Tiny Story [eldan2023tinystories] | 172.33 | 7073.81 | 10477 | 1000 | - | ||||||||
-| user review | IMDB [imdb_reviews_dataset] | 194.06 | 8761.54 | 30569 | 1000 | IMDB License | ||||||||
-| user review | yelp [asghar2016yelp] | 128.16 | 5542.79 | 17244 | 1000 | Yelp License | ||||||||
-| essay | asap-aes [asap_aes_dataset] | 199.43 | 8734.33 | 15142 | 1000 | - | ||||||||
-| overall | - | 171.3113 | 7749.073 | 111662 | 8000 | - |
-Appendix B Experimental Details
+7.4 IVTG
+To further assess the ability of our , we construct a new dataset named IVTG, which is entirely synthesized by the LLM.
+Specifically, we leverage the DeepSeek [14] API to generate the textual content.
+Similar to the original IVT dataset, we divide IVTG into three subsets based on the length of the secret text.
+8 Experimental Details
 To explore the feasibility and performance of sentence-to-image steganography powered by large language models, we propose three variants of our framework, namely -Qwen2.5-0.5B, -Llama3.2-1B, and -MiniCPM-1B.
@@
 This enables a thorough evaluation of how different LLMs perform under the semantic steganography task.
-B.1 Implementation Details
+For all baseline methods, we follow the training settings described in their original papers to ensure a fair comparison.
+8.1 Training Dataset of
+Text Dataset.
+We use the English subset of the WanJuan [25] Text Dataset as our text dataset.
+WanJuan 1.0 Text Dataset is composed of cleaned pre-training corpora from different sources such as web pages, encyclopedias, books, patents, textbooks, and exam questions.
+The total amount of data exceeds 500 million documents, and the data size exceeds 1TB.
+We selected a subset of the English dataset (4 GB).
+For texts that spanned multiple lines, we removed ββ, ββ and extra spaces to merge them into single lines.
+We then extracted texts with fewer than 1,000 words, resulting in 3,537,390 samples.
+Finally, we randomly selected 320,000 samples to construct the training set.
+The distribution of text lengths in the training set is shown in Fig.Β 7.
+Image Dataset. We use the COCO Training Dataset [41] training dataset as our cover image dataset.
+8.2 Implementation Details
 LoRA Config.
@@
 -Qwen2.5-0.5B.
--Qwen2.5-0.5B is built upon the Qwen2.5-0.5B backbone [qwen2025qwen25technicalreport] and fine-tuned using the LoRA config mentioned above.
+-Qwen2.5-0.5B is built upon the Qwen2.5-0.5B backbone [55] and fine-tuned using the LoRA config mentioned above.
 The training strategy is detailed in Section 4.3.
@@
 -Llama3.2-1B.
--Llama3.2-1B is built upon the Llama3.2-1B backbone [grattafiori2024llama] and fine-tuned using the LoRA config mentioned above.
+-Llama3.2-1B is built upon the Llama3.2-1B backbone [21] and fine-tuned using the LoRA config mentioned above.
 The training strategy is detailed in Section 4.3.
@@
 -MiniCPM-1B.
--MiniCPM-1B is built upon the MiniCPM-1B backbone [hu2024minicpm] and fine-tuned using the LoRA config mentioned above.
+-MiniCPM-1B is built upon the MiniCPM-1B backbone [28] and fine-tuned using the LoRA config mentioned above.
 The training strategy is detailed in Section 4.3.
@@
 -Gemma3-1B.
--Gemma3-1B is built upon the Gemma3-1B backbone [team2025gemma] and fine-tuned using the LoRA config mentioned above.
+-Gemma3-1B is built upon the Gemma3-1B backbone [61] and fine-tuned using the LoRA config mentioned above.
 The training strategy is detailed in Section 4.3.
@@
 StegaStamp.
-We retrain StegaStamp [tancik2020stegastamp] with extended capacity (from 100 to 512 bits) on cover images with 256Γ256 resolution using the Flickr2K dataset.
+We retrain StegaStamp [60] with extended capacity (from 100 bits to 512 bits) on cover images with resolution using the Flickr2K dataset.
 The model is optimized with AdamW with learning rate set to 2e-4 for 15,000 iterations with batch size set to 16.
-While achieving stable convergence, we observe fundamental limitations: the 512-bit encoding yields only ~70% bit accuracy, resulting terrible performance in IVT.
+While achieving stable convergence, we observe fundamental limitations: the 512-bit encoding yields only 70% bit accuracy, resulting terrible performance in IVT.
 This suggests current spatial-domain watermarking architectures struggle with high-capacity payloads with primitive representation of secret message such as UTF codes.
+More details can be found in Sec.Β 12.2.
 LanNet.
-We retrain LanNet [lan2023robust] on cover images with 256Γ256 resolution using the DIV2K training dataset.
-The model is optimized with AdamW with learning rate set to 2e-4 for 2,000 epochs with batch size set to 16.
-B.2 Details about Training Dataset
-To construct the training dataset for the aforementioned models, we curated a diverse collection of textual data from multiple sources.
-Specifically, we selected 14,000 highlights from the CNN/DailyMail dataset [see-etal-2017-get], 5,000 news headlines from the PENS dataset [ao-etal-2021-pens], 5,000 scientific titles from the ArXiv-10 dataset [farhangi2022protoformer], and 4,000 story paragraphs from the CNN Story corpus [cnn_stories_dataset].
-This composition ensures a broad range of semantic structures and topics, facilitating robust model training across different text granularities.
-Importantly, we ensure that there is no overlap between the training set and the proposed IVT benchmark datasets, thereby preventing data leakage and preserving the integrity of the evaluation.
-For more details about IVT-L, please refer to Table 7.
-| Category | Source |
-| news headlines | PENS [ao-etal-2021-pens] | 21.6 | 1296.3 | 5000 | ||||||
-| paper title | arxiv-10 [farhangi2022protoformer] | 25.1 | 1402.0 | 5000 | ||||||
-| news highline | CNN/DailyMail [see-etal-2017-get] | 55.8 | 2826.4 | 14000 | ||||||
-| new story | CNN Story [cnn_stories_dataset] | 362.6 | 11024.6 | 4000 |
-Appendix C Steganalysis
-C.1 Statistical Steganalysis
-We use StegExpose [Boehm2014StegExposeA] to measure the anti-steganalysis ability.
-The receiver operating characteristic (ROC) curve in FigureΒ 8 shows that has high security and can fool the StegExpose.
-Appendix D VLMs in
-In the previous experiments, all models were based on pure text-only large language models.
-To further explore the applicability of vision-language models (VLMs) to the sentence-to-image steganography task, we conducted additional experiments using Qwen2-vl [Qwen2-VL] as a representative VLM.
-However, the results showed that Qwen2-vl failed to fit the training data.
-We hypothesize that this underperformance stems from the fundamental difference between visual understanding and steganography.
-While VLMs are primarily optimized for visual tasks, which focusing on understanding objects, scenes, and spatial relationships within images.
-However, steganography emphasizes decoding secret messages by focusing on subtle, imperceptible signals embedded within images.
-Unlike typical vision-language alignment tasks, steganography does not require high-level visual semantics but rather the capacity to detect and interpret fine-grained patterns that are deliberately encoded.
-This highlights the need for models specifically tailored to the semantic embedding and recovery processes intrinsic to image-based steganography.
-In experiments, the -Qwen2-vl-2B model can output the and tokens, but the recovery information is totally different from the ground truth.
-That may because the knowledge in the vision encoder seriously hindering the language model to get exact secret information.
-Appendix E Experiments on Generative Message
-To further assess the ability of our , we construct a new dataset named IVTG, which is entirely synthesized by the LLM.
-Specifically, we leverage the DeepSeek [deepseek2024] API to generate the textual content, as illustrated in Appendix Section F.
-Similar to the original IVT dataset, we divide IVTG into three subsets based on the length of the secret text.
-As shown in Table 8, our model achieves comparable performance on IVTG to its performance on the original IVT dataset.
-This consistency demonstrates the modelβs ability to generalize to data generated by large language models, further validating the effectiveness of our approach.
-Appendix F Codes for Collecting IVTG
+We retrain LanNet [36] on cover images with resolution using the DIV2K training dataset.
+The model is optimized with AdamW with learning rate set to 2e-4 for 1,000 epochs with batch size set to 12.
+FPGP.
+We retrain FPGP [78] on cover images with resolution using the DIV2K training dataset.
+The model is optimized with AdamW with learning rate set to 2e-4 for 1,000 epochs with batch size set to 12.
+CRMark.
+We retrain CRMark [10] on cover images with resolution using the DIV2K training dataset.
+The model is optimized with AdamW with learning rate set to 1e-4 for 1000 epochs with batch size set to 12.
+| Methods | -Q | DwtDct | StegaStamp | LanNet | FPGP | CRMark |
+| K-L | 0.0015 | 0.082 | 0.072 | 0.0024 | 0.0017 | 0.0019 |
+9 Statistical Steganalysis
+We use StegExpose [6] to measure the anti-steganalysis ability. The receiver operating characteristic (ROC) curve in FigureΒ 9 shows that has high security and can fool the StegExpose.
+K-L divergence provides a convenient means of gauging how easy it is to discriminate between cover and stego.
+Typical cover data, however, are not i.i.d.
+(independent and identically distributed).
+For example, both pixels and audio samples are known to be highly correlated [58].
+Therefore, KL divergence is insufficient as a standalone criterion for evaluating steganographic security [58].
+Nevertheless, we compute the K-L divergence between the stego images generated by and the original cover images.
+As shown in Tab.Β 7, all steganographic models achieve very small KL divergence values, with getting the best result.
 | IVTG-S | IVTG-M | IVTG-L | ||||||||||||||||
@@
 […]
-Appendix G Limitations
-While demonstrates strong performance in encoding and recovering semantically rich text, it exhibits limitations when handling non-semantic or low-semantic content, such as random digit strings, binary sequences, or nonsensical phrases.
-This limitation contrasts with some watermarking-based frameworks, which are designed to handle raw bit-level information and therefore remain agnostic to the semantic nature of the payload.
-In contrast, leverages a language modeling backbone that inherently assumes the presence of syntactic and semantic structures in the secret message.
-Appendix H Broader Impacts
+10 Supplementary Experiments
+10.1 VLMs in
+In the previous experiments, all models were based on pure text-only large language models.
+To further explore the applicability of vision-language models (VLMs) to the sentence-to-image steganography task, we conducted additional experiments using Qwen2-vl [64] as a representative VLM.
+However, the results showed that Qwen2-vl failed to fit the training data.
+Fig.Β 8 shows a failure case of -Qwen2-vl-2B on IVT-L.
+The model can output the and tokens.
+However, the recovered text is entirely different from the ground-truth message.
+We hypothesize that this underperformance stems from the fundamental difference between visual understanding and steganography.
+While VLMs are primarily optimized for visual tasks, which focusing on understanding objects, scenes, and spatial relationships within images.
+However, steganography emphasizes decoding secret messages by focusing on imperceptible signals embedded within images.
+Unlike typical vision-language alignment tasks, steganography does not require high-level visual semantics but rather the capacity to detect and interpret fine-grained patterns that are deliberately encoded.
+This highlights the need for models specifically tailored to the semantic embedding and recovery processes intrinsic to image-based steganography.
+To validate our hypothesis, we conducted an ablation study by removing the visual encoder of Qwen2-vl.
+As shown in Tab.Β 9, under this setting, the Qwen2-vl works well on IVT-L, suggesting that its original visual encoder may in fact hinder its ability to capture low-level steganographic signals.
+| Methods | WER | BLEU-4 | ROUGE | BERT-S | PSNR | SSIM |
+| -Qwen2-vl-wo_VE | 0.982 | 0.012 | 0.035 | 0.106 | 33.51 | 0.827 |
+| -Qwen2-vl-w_VE | 0.152 | 0.841 | 0.832 | 0.890 | 32.16 | 0.812 |
+| Methods | WER | BLEU-4 | ROUGE | BERT-S | PSNR | SSIM |
+| -Qwen2.5-0.5B-8bit | 0.201 | 0.804 | 0.882 | 0.907 | 40.05 | 0.946 |
+10.2 Model Quantization
+We use a lightweight language model (), which runs efficiently on edge devices such as mobile phones.
+After applying 8-bit quantization, its memory usage of the -Qwen2.5-0.5B is below 200 MB, making the entire framework both practical and scalable.
+As shown in the table below, the 8-bit quantized version of -Qwen2.5-0.5B processes the IVT-L dataset effectively without any quantization-aware training, with GPU memory usage of less than 2 GB during inference.
+10.3 Experiments on Generative Message
+To further assess the ability of our , we construct a new dataset named IVTG, which is entirely synthesized by the LLM.
+Specifically, we leverage the DeepSeek [14] API to generate the textual content.
+Similar to the original IVT dataset, we divide IVTG into three subsets based on the length of the secret text.
+As shown in Table 8, our model achieves comparable performance on IVTG to its performance on the original IVT dataset.
+This consistency demonstrates the modelβs ability to generalize to data generated by large language models, further validating the effectiveness of our approach.
+10.4 Experiments on Non-semantic Message
+Under the definition of semantic steganography, secret messages are assumed to be natural language information.
+All of our experiments follow this assumption.
+However, not all text carries semantic content.
+For instance, a random character string or a sequence of bits conveys no meaningful information, and transmitting such data offers limited practical utility.
+Nevertheless, some forms of non-semantic text do appear in real-world scenarios, such as URLs or cryptographic keys.
+To evaluate these cases, we tested on several representative examples and visualized the results.
+As shown in Fig.Β 10, -Qwen2.5-0.5B maintains strong performance even in these settings.
+| Methods | WER | BLEU-4 | ROUGE | BERT-S |
+| Original Qwen2.5-0.5B | 1.000 | 0.000 | 0.003 | 0.336 |
+| Different Qwen2.5-0.5B | 1.000 | 0.000 | 0.004 | 0.324 |
+10.5 Decoding with the Original LLM
+We evaluate the security of under a stronger and more practical adversarial assumption: the adversary successfully identifies the exact LLM used in our framework.
+Specifically, we conduct experiments on -Qwen2.5-0.5B and assume that the adversary knows the concrete LLM and has even obtained the Patch-to-Token MLP .
+In this setting, we generate stego images using the fine-tuned Qwen2.5-0.5B, while the adversary attempts to decode the hidden messages using the original Qwen2.5-0.5B.
+All prompts and decoding configurations remain unchanged.
+As illustrated in Tab.Β 11 (Original Qwen2.5-0.5B), the decoding results are extremely poor: the recovered text contains little to no semantic correspondence to the ground truth.
+This observation confirms that, even when the adversary correctly identifies the underlying LLM and gains access to , the hidden information cannot be decoded without the fine-tuned decoding alignment.
+In conclusion, remains secure under this strong known-model assumption if the weight of the LLM is not leaked.
+10.6 Decoding Using Cross-Instance LLM
+We further assess the security of by examining whether independently trained model instances are mutually decodable.
+Using identical training configurations but different random seeds, we obtain two variants of -Qwen2.5-0.5B, denoted as -A and -B.
+In this experiment, we use -A to generate stego image , while -B is used for decoding information from stego image.
+Despite sharing the same architecture, training hyperparameters, and optimization procedure, -B fails to recover meaningful messages from stego images generated by -A, as shown in the Tab.Β 11 (Different Qwen2.5-0.5B).
+These findings indicate that the steganographic latent alignment learned during fine-tuning is instance-specific and not transferable across independently trained models, further demonstrating the robustness of to cross-instance decoding attempts.
+11 More Details about Ablation Study
+11.1 Impact of Mask Strategy
+In decoding process of the first stage, we mask some input tokens to encourage LLM distribute the information in the whole cover image.
+As shown in FigureΒ 11, the SMEs are evenly distributed after training with the mask strategy, and significant improvements in recovery quality are observed in TableΒ 3.
+12 Analysis of Existing Methods
+12.1 Choice of the Baselines
+To ensure a fair and reasonable comparison, we select several representative methods from the field of image steganography as our baselines.
+Our chosen baselines include both traditional methods (DwtDct) and deep learning-based models (StegaStamp, LanNet, FPGP, CRMark).
+In the original paper of StegaStamp, the authors categorize their method as steganography.
+However, following the terminology adopted in prior literature, we classify it as a watermarking method, as it takes robustness into considerations.
+However, in our experiments, we remove the robustness constraint to align them with the definition of steganography.
+All deep learning-based baselines are retrained for a fair comparison.
+12.2 Why Existing Model Fail?
+As illustrated in the Fig.Β 12, we show the architectural details of existing methods.
+To facilitate further analysis, we categorize existing models into two types based on how they process the secret message: vector-base models [60, 83, 68] and image-based models [36, 78, 10].
+Vector-based methods employ MLPs to transform bitstreams into vector representation.
+During the embedding process, an embedding MLP first converts the bit sequence into a latent vector, which is then reshaped to match the spatial resolution of the cover image.
+This latent representation is subsequently fused with the cover image through an image encoder to produce the stego image.
+For message recovery, the stego image is processed by an image decoder to extract feature maps, which are flattened and passed through a decoding MLP to reconstruct the original binary message.
+In contrast, image-based methods omit the use of MLPs.
+Instead, the raw bit sequence is directly reshaped into a two-dimensional (2D) map and up-sampled (up-sample is optional) to align with the resolution of cover image.
+Then, the 2D representation of bitstreams will concat with cover image to produce stego image by image encoder.
+The stego image is then decoded by an image decoder to recover the 2D secret map, which is average pooled (pooling is optional) and flattened to reconstruct the final message.
+All of these two types of approaches encode textual information into bitstreams using UTF encoding for embedding and extraction.
+While this design is straightforward, it fails to preserve the semantic characteristics of the original text.
+In our experiments reported in the main paper, we found that existing models exhibit a notable limitation in steganographic capacity.
+Although some of these methods were originally evaluated under watermarking task settings, where a trade-off between robustness and capacity is necessary, our additional experiments in Sec.Β 12.4 suggest that robustness is not the primary factor causing this limitation.
+Instead, the architectural design fundamentally constrains the upper bound of steganographic capability.
+This structural limitation has been largely overlooked in previous studies, overshadowed by concerns about robustness.
+In summary, existing methods face two main limitations, as described below.
+Limitations of Binary Representation.
+Existing methods typically employ UTF-based encoding for secret messages.
+Specifically, the textual content is first converted into a binary sequence through UTF encoding and then processed by subsequent modules.
+However, this binary representation introduces several inherent limitations.
+(1) Low information density.
+UTF encoding is a manually designed one-to-one mapping that lacks linguistic structures such as words, morphemes, or syntax.
+Consequently, it provides no inductive bias, hindering the modelβs ability to learn meaningful representations in subsequent procedures.
+(2) Excessive sequence length. The UTF encoding scheme represents each character with 8 bits, leading to excessively long binary sequences that are inefficient and cumbersome to process.
+(3) Limited capacity and scalability.
+Both types of methods require reshaping the binary sequence to align with the spatial resolution of the image.
+Consequently, a maximum bit length must be predefined: shorter sequences are padded, whereas longer ones are truncated.
+As illustrated in Fig.Β 12, in vector-based methods using an MLP, both the input dimension N and output dimension M are fixed.
+Similarly, in image-based methods where the binary sequence is directly concatenated with image features, the reshaped sequence must match the pixel resolution.
+Therefore, existing models can only process fixed-length bit sequences, it is a rigid constraint that also imposes a fixed resolution on the cover image.
+Limitations of MLP.
+Once the secret message is converted into a binary sequence, it is typically processed by a multi-layer perceptron (MLP) to obtain a latent representation.
+However, this operation introduces several critical issues:
+(1) Poor generalization.
+As a fully connected network, the MLP lacks the ability to capture local or global dependencies within the input sequence.
+Moreover, binary encoding is inherently structureless, which severely reduces representational efficiency.
+Since UTF-based binary representations carry no linguistic inductive bias, the MLP struggles to learn meaningful or generalizable patterns.
+(2) Limited capacity and flexibility.
+Because the MLPβs input and output dimensions are fixed, it can only handle binary sequences of a specific length.
+This rigid design constrains the modelβs capacity and prevents it from adapting to messages of variable length.
+(3) Parameter explosion.
+When the binary sequence is long, the MLP must process extremely high-dimensional inputs.
+For example, with an input length of 4,096 bits and a hidden dimension of 512, the first layer alone contains more than two million parameters, rendering the model inefficient to train.
+To validate our conclusion that the architectural design of existing methods fundamentally limits the capabilities of current steganographic techniques, we conduct comprehensive experiments in the following section.
+The results indicate that the observed capacity bottleneck stems from the model architecture rather than from robustness considerations.
+Inverse Wooden Barrel Effect.
+Traditional steganographic models are typically trained for a fixed maximum message capacity.
+Under this setting, the model must accommodate the longest possible message, even though real-world communications rarely reach this limit.
+As a result, the architecture is forced to scale its internal capacity, likes constructing a barrel tall enough to hold the deepest possible fill.
+In practice, however, only a small portion of this capacity is used, leaving most of it wasted.
+We term this mismatch the Inverse Barrel Effect.
+In contrast to the classical barrel effect, where the shortest stave limits the total volume, here the longest stave determines the overall design.
+The model must be overparameterized to satisfy the largest capacity requirement, resulting in redundant computation, and potentially degraded performance when embedding short messages.
+| Capacity | PSNR | SSIM |
+| 256 bit | 45.91 | 0.9792 | 1.0000 | 1.0000 | ||||
+| 512 bit | 43.40 | 0.9568 | 0.7597 | 0.0553 | ||||
+| 768 bit | 40.48 | 0.9103 | 0.6910 | 0.0222 | ||||
+| 1024 bit | 50.12 | 0.9999 | 0.5039 | 0.0021 |
+| Capacity | PSNR | SSIM |
+| 0.015 bpp (1024 bits) | 45.92 | 0.9892 | 1.0000 | 1.0000 | ||||
+| 0.0625 bpp (4096 bits) | 37.02 | 0.9201 | 1.0000 | 1.0000 | ||||
+| 0.25 bpp (16384 bits) | 30.10 | 0.8115 | 0.5516 | 0.0035 | ||||
+| 1 bpp (65536 bits) | 25.18 | 0.6602 | 0.5208 | 0.0038 |
+12.3 Capacity Analysis of Existing Models.
+To validate the above analysis, we evaluated the capacity limits of existing models after removing robustness constraints.
+We selected two representative methods, StegaStamp (vector-based method) and CRMark (image-based method).
+For StegaStamp, we expanded its capacity by adjusting the embedding MLP to support 256, 512, 768, and 1024 bits.
+For CRMark, we tested the model under different payload rates of 0.015625, 0.0625, 0.25, and 1 bits per pixel (bbp).
+The resolution of cover image is fixed to .
+| Attack | Methods | IVT-S | IVT-M | IVT-L | |||||||||||||||
+| Secret/Recovery | Cover/Stego | Secret/Recovery | Cover/Stego | Secret/Recovery | Cover/Stego | ||||||||||||||
+| WER | BLEU | ROUGE | BERT-S | PSNR | SSIM | WER | BLEU | ROUGE | BERT-S | PSNR | SSIM | WER | BLEU | ROUGE | BERT-S | PSNR | SSIM | ||
+| Gaussian Blur | StegaStamp | 0.402 | 0.231 | 0.125 | 0.332 | - | - | - | - | - | - | - | - | - | - | - | - | - | - |
+| DwtDct | 1.000 | 0.000 | 0.000 | 0.292 | - | - | - | - | - | - | - | - | - | - | - | - | - | - | |
+| FPGP | 0.134 | 0.602 | 0.935 | 0.922 | - | - | 0.821 | 0.485 | 0.565 | 0.491 | - | - | - | - | - | - | - | - | |
+| LanNet | 0.010 | 0.783 | 0.927 | 0.989 | - | - | 0.251 | 0.682 | 0.782 | 0.815 | - | - | 0.910 | 0.056 | 0.142 | 0.382 | - | - | |
+| CRMark | 0.012 | 0.735 | 0.952 | 0.962 | - | - | 0.173 | 0.752 | 0.802 | 0.863 | - | - | 0.859 | 0.102 | 0.147 | 0.420 | - | - | |
+| -Q | 0.042 | 0.842 | 0.883 | 0.965 | - | - | 0.042 | 0.961 | 0.978 | 0.975 | - | - | 0.081 | 0.912 | 0.972 | 0.965 | - | - | |
+| Brightness | StegaStamp | 0.443 | 0.252 | 0.251 | 0.389 | - | - | - | - | - | - | - | - | - | - | - | - | - | - |
+| DwtDct | 1.000 | 0.000 | 0.000 | 0.304 | - | - | - | - | - | - | - | - | - | - | - | - | - | - | |
+| FPGP | 0.141 | 0.721 | 0.893 | 0.904 | - | - | 0.747 | 0.434 | 0.552 | 0.454 | - | - | - | - | - | - | - | - | |
+| LanNet | 0.031 | 0.715 | 0.982 | 0.927 | - | - | 0.251 | 0.621 | 0.812 | 0.825 | - | - | 0.914 | 0.062 | 0.120 | 0.362 | - | - | |
+| CRMark | 0.021 | 0.752 | 0.994 | 0.995 | - | - | 0.157 | 0.782 | 0.811 | 0.852 | - | - | 0.876 | 0.155 | 0.176 | 0.402 | - | - | |
+| -Q | 0.052 | 0.821 | 0.893 | 0.978 | - | - | 0.062 | 0.934 | 0.981 | 0.961 | - | - | 0.073 | 0.921 | 0.951 | 0.964 | - | - | |
+| S&P Noise | StegaStamp | 0.408 | 0.226 | 0.125 | 0.341 | - | - | - | - | - | - | - | - | - | - | - | - | - | - |
+| DwtDct | 1.000 | 0.000 | 0.000 | 0.284 | - | - | - | - | - | - | - | - | - | - | - | - | - | - | |
+| FPGP | 0.131 | 0.762 | 0.911 | 0.936 | - | - | 0.801 | 0.512 | 0.571 | 0.497 | - | - | - | - | - | - | - | - | |
+| LanNet | 0.020 | 0.715 | 0.985 | 0.976 | - | - | 0.216 | 0.714 | 0.828 | 0.835 | - | - | 0.952 | 0.027 | 0.142 | 0.290 | - | - | |
+| CRMark | 0.052 | 0.752 | 0.972 | 0.988 | - | - | 0.200 | 0.704 | 0.831 | 0.862 | - | - | 0.893 | 0.135 | 0.127 | 0.395 | - | - | |
+| -Q | 0.047 | 0.876 | 0.902 | 0.954 | - | - | 0.038 | 0.937 | 0.978 | 0.966 | - | - | 0.072 | 0.932 | 0.957 | 0.974 | - | - | |
+| Gaussian Noise | StegaStamp | 0.461 | 0.272 | 0.164 | 0.333 | - | - | - | - | - | - | - | - | - | - | - | - | - | - |
+| DwtDct | 1.000 | 0.000 | 0.000 | 0.294 | - | - | - | - | - | - | - | - | - | - | - | - | - | - | |
+| FPGP | 0.106 | 0.751 | 0.918 | 0.933 | - | - | 0.690 | 0.531 | 0.572 | 0.555 | - | - | - | - | - | - | - | - | |
+| LanNet | 0.021 | 0.751 | 0.961 | 0.986 | - | - | 0.201 | 0.702 | 0.856 | 0.815 | - | - | 0.921 | 0.022 | 0.104 | 0.346 | - | - | |
+| CRMark | 0.019 | 0.812 | 0.974 | 0.996 | - | - | 0.192 | 0.742 | 0.886 | 0.874 | - | - | 0.896 | 0.102 | 0.125 | 0.402 | - | - | |
+| -Q | 0.041 | 0.872 | 0.884 | 0.962 | - | - | 0.064 | 0.967 | 0.977 | 0.961 | - | - | 0.062 | 0.942 | 0.961 | 0.956 | - | - | |
+| Rescale | StegaStamp | 0.441 | 0.236 | 0.126 | 0.362 | - | - | - | - | - | - | - | - | - | - | - | - | - | - |
+| DwtDct | 1.000 | 0.000 | 0.000 | 0.306 | - | - | - | - | - | - | - | - | - | - | - | - | - | - | |
+| FPGP | 0.132 | 0.682 | 0.914 | 0.928 | - | - | 0.731 | 0.501 | 0.511 | 0.586 | - | - | - | - | - | - | - | - | |
+| LanNet | 0.021 | 0.751 | 0.970 | 0.984 | - | - | 0.212 | 0.731 | 0.825 | 0.864 | - | - | 0.904 | 0.078 | 0.103 | 0.368 | - | - | |
+| CRMark | 0.017 | 0.852 | 0.980 | 0.995 | - | - | 0.214 | 0.841 | 0.892 | 0.901 | - | - | 0.845 | 0.121 | 0.124 | 0.341 | - | - | |
+| -Q | 0.043 | 0.882 | 0.895 | 0.952 | - | - | 0.042 | 0.961 | 0.978 | 0.970 | - | - | 0.068 | 0.926 | 0.944 | 0.967 | - | - | |
+| JPEG | StegaStamp | 0.431 | 0.226 | 0.151 | 0.362 | - | - | - | - | - | - | - | - | - | - | - | - | - | - |
+| DwtDct | 1.000 | 0.000 | 0.000 | 0.304 | - | - | - | - | - | - | - | - | - | - | - | - | - | - | |
+| FPGP | 0.169 | 0.732 | 0.911 | 0.931 | - | - | 0.821 | 0.471 | 0.482 | 0.561 | - | - | - | - | - | - | - | - | |
+| LanNet | 0.041 | 0.741 | 0.952 | 0.984 | - | - | 0.201 | 0.737 | 0.792 | 0.862 | - | - | 0.942 | 0.020 | 0.094 | 0.401 | - | - | |
+| CRMark | 0.025 | 0.851 | 0.962 | 0.982 | - | - | 0.194 | 0.804 | 0.916 | 0.921 | - | - | 0.832 | 0.102 | 0.156 | 0.421 | - | - | |
+| -Q | 0.051 | 0.861 | 0.901 | 0.951 | - | - | 0.041 | 0.958 | 0.983 | 0.959 | - | - | 0.072 | 0.952 | 0.962 | 0.982 | - | - | |
+| Average | StegaStamp | 0.431 | 0.280 | 0.157 | 0.411 | 33.6 | 0.807 | - | - | - | - | - | - | - | - | - | - | - | - |
+| DwtDct | 1.000 | 0.000 | 0.000 | 0.297 | 31.8 | 0.898 | - | - | - | - | - | - | - | - | - | - | - | - | |
+| FPGP | 0.135 | 0.708 | 0.913 | 0.925 | 40.3 | 0.857 | 0.768 | 0.489 | 0.542 | 0.524 | 35.6 | 0.874 | - | - | - | - | - | - | |
+| LanNet | 0.024 | 0.742 | 0.962 | 0.974 | 40.4 | 0.861 | 0.222 | 0.698 | 0.816 | 0.836 | 35.3 | 0.892 | 0.923 | 0.044 | 0.118 | 0.358 | 29.6 | 0.795 | |
+| CRMark | 0.024 | 0.792 | 0.972 | 0.986 | 42.7 | 0.895 | 0.188 | 0.770 | 0.856 | 0.878 | 37.2 | 0.910 | 0.866 | 0.120 | 0.142 | 0.396 | 30.1 | 0.811 | |
+| -Q | 0.046 | 0.859 | 0.893 | 0.960 | 39.0 | 0.9718 | 0.048 | 0.953 | 0.979 | 0.965 | 35.41 | 0.9519 | 0.007 | 0.930 | 0.957 | 0.968 | 30.4 | 0.8345 | |
+Our findings are consistent with recent results reported by Meta [52].
+For StegaStamp, when the capacity is increased to 1024 bits, the model exhibits clear mode collapse: the decoder fails to recover any meaningful secret information, and the encoder also stops functioning properly.
+Similarly, for CRMark, its decoding performance degrades substantially when the payload reaches 0.25 bpp.
+Notably, this decrease in decoding accuracy is accompanied by a decline in quality of stego image, which may stem from the inherent invertibility constraints of the INN architecture.
+Consistent with these observations [52], researchers reported that VideoSeal fails to encode more than 1024 bits when evaluated solely under PSNR constraints.
+They similarly attempted to scale VideoSealβs capacity and, after considerable effort, introduced ChunkySeal, which increases the payload from 256 bits to 1024 bits.
+However, this improvement comes at a substantial cost: ChunkySeal is several orders of magnitude larger than VideoSeal.
+Its encoder contains 1B parameters and its decoder has 773M parameters, far exceeding those of VideoSeal (11M Encoder and 33M Decoder) and even surpassing our own model.
+These experiments validate our analysis: binary representations impose inherent constraints on model design, thereby limiting steganographic capacity.
+As shown in the subsequent robustness experiments, consistently maintains strong performance across diverse attack scenarios, further demonstrating the effectiveness of our proposed pipeline.
+However, a more detailed investigation into why existing models struggle to scale their capacity can be left to future work.
+12.4 Robustness Study.
+To evaluate the robustness of , we trained the model with simulated attacks follow previous works. However, it is worth noting that robustness is not a standard requirement in steganography tasks.
+Experimental Settings. We introduced several common distortions, including gaussian noise, salt-and-pepper noise (S&P noise), gaussian blur, rescaling, bright adjustment, and JPEG compression.
+For Gaussian noise, we injected zero-mean noise with standard deviations to simulate varying intensity levels (pixel value range is ).
+For salt-and-pepper noise, the corruption ratio was set to .
+For Gaussian blur, we applied kernels of size with a fixed standard deviation of 1.0.
+For rescaling, images were resized to of their original resolution and then sampled using bilinear interpolation.
+For brightness adjustment, pixel intensities were scaled by a random factor within the range .
+Finally, JPEG compression with quality factors was used to assess compression robustness, we employ differentiable-JPEG to simulate the compression process.
+We follow the same training configuration as in the main experiments, with a single modification in the decoding stage: before feeding the stego image to the LLM for recovery, we randomly apply a combination of one to three attacks to the stego image.
+Experimental Results.
+As shown in Tab.Β 14, remains robust under all attack settings.
+It preserves strong decoding performance without any loss of capacity.
+In contrast, the baseline models degrade substantially once attacks are introduced, exhibiting significant drops in recovery accuracy across all settings and highlighting the advantages of our pipeline.
+This experiment demonstrates that once perturbation simulation is integrated into training, can reliably perform the watermarking task while maintaining superior capacity.
+13 Discussion
+13.1 Why We Need Semantic Steganography
+Regardless of whether information is represented as tokens or bitstreams, the ultimate goal of steganography is to secretly transmit human-interpretable information.
+Semantic steganography redefines what can be hidden, elevating steganography to a higher-level abstraction.
+At this level, the focus is no longer on whether individual bits are correctly decoded, instead, it emphasizes whether the intended information is successfully and secretly conveyed.
+In future research, the philosophy of semantic steganography can be naturally extended to multi-modal steganography, digital watermarking, copyright protection, and other related applications.
+In summary, semantic steganography opens up a new direction for the development of steganographic models.
+13.2 Why Not LSB
+Over the past two decades, many traditional steganographic methods have been proposed, such as Least Significant Bit (LSB) substitution and transform-domain techniques based on Discrete Cosine Transform (DCT) or Discrete Wavelet Transform (DWT).
+These approaches are generally categorized as cover-modification steganography [19].
+However, their underlying mechanism is relatively simple.
+They embed secret information by directly replacing certain discrete units of the cover, such as the least significant bits of pixels or frequency coefficients.
+Therefore, we refer to them as substitution-based steganography.
+Despite their simplicity, this class of methods suffers from several critical limitations:
+β’
+Low security: These methods are hand-crafted and tend to introduce noticeable statistical artifacts.
+Over the past decades, a large number of steganalysis techniques have been developed to detect such artifacts, rendering these methods essentially insecure in practice.
+β’
+Poor scalability: The substitution operation is overly simple and cannot be easily extended to other tasks such as watermarking.
+In addition, this operation is non-differentiable, making it difficult to integrate with modern deep learning models.
+In contrast, our approach does not suffer from the above issues.
+Although it also belongs to cover-modification steganography, leverages large language models (LLMs) to automatically generate secret message embeddings (SMEs), eliminating the need for manual design.
+The generated embeddings exhibit more complex and less predictable structures, which provide stronger resistance against classical steganalysis.
+To evaluate this property, we applied the RS analysis to both the traditional LSB method and .
+As shown in Tab.Β 15, easily withstands this statistical analysis, while the LSB method can be detected with high confidence.
+| Method | -Qwen2.5-0.5B | LSB | Clean |
+| RS Analysis | 0.102 | 0.254 | 0.098 |
+13.3 Broader Impacts
 Semantic steganography introduces a novel paradigm that leverages carriers to conceal the structure and semantics of natural language, enabling new possibilities beyond traditional bit-level encoding schemes.
@@
 Embedding meaningful language into other media formats presents opportunities for developing cross-modal steganographic systems capable of high-capacity, human-aligned, and content-aware data hiding.
+Such techniques may benefit fields such as digital rights management, covert communication, interactive storytelling, and information watermarking in generative media.
 We hope that this work not only advances semantic steganography from a technical perspective but also encourages discussion around its responsible use, ethics, and governance frameworks in future applications.
-Appendix I Additional Results
-NeurIPS Paper Checklist
-1.
-Claims
-Question: Do the main claims made in the abstract and introduction accurately reflect the paperβs contributions and scope?
-Answer: [Yes]
-Justification: The abstract and introduction clearly state the proposed concept of semantic steganography, introduce the sentence-to-image task, and present the , accurately reflecting the paperβs contributions and scope.
-Guidelines:
-β’
-The answer NA means that the abstract and introduction do not include the claims made in the paper.
-β’
-The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations.
-A No or NA answer to this question will not be perceived well by the reviewers.
-β’
-The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
-β’
-It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
-β’
-2.
-Limitations
-Question: Does the paper discuss the limitations of the work performed by the authors?
-Answer: [Yes]
-Justification: We discusses its limitations in the appendix G, acknowledging areas for future improvement and potential constraints of the proposed approach.
-Guidelines:
-β’
-The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
-β’
-The authors are encouraged to create a separate "Limitations" section in their paper.
-β’
-The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally).
-The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
-β’
-The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs.
-In general, empirical results often depend on implicit assumptions, which should be articulated.
-β’
-The authors should reflect on the factors that influence the performance of the approach.
-For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting.
-Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
-β’
-The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
-β’
-If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
-β’
-While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that arenβt acknowledged in the paper.
-The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community.
-Reviewers will be specifically instructed to not penalize honesty concerning limitations.
-β’
-3.
-Theory assumptions and proofs
-Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
-Answer: [No]
-Justification: No theoretical result
-Guidelines:
-β’
-The answer NA means that the paper does not include theoretical results.
-β’
-All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
-β’
-All assumptions should be clearly stated or referenced in the statement of any theorems.
-β’
-The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
-β’
-Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
-β’
-Theorems and Lemmas that the proof relies upon should be properly referenced.
-β’
-4.
-Experimental result reproducibility
-Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
-Answer: [Yes]
-Guidelines:
-β’
-The answer NA means that the paper does not include experiments.
-β’
-If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
-β’
-If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
-β’
-Depending on the contribution, reproducibility can be accomplished in various ways.
-For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model.
-In general.
-releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
-β’
-While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution.
-For example
-If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
-If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
-If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
-We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility.
-In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
-β’
-5.
-Open access to data and code
-Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
-Answer: [Yes]
-Justification: The IVT dataset is included in supplemental material, the code of will be released after the paper being accepted.
-Guidelines:
-β’
-The answer NA means that paper does not include experiments requiring code.
-β’
-Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
-β’
-While we encourage the release of code and data, we understand that this might not be possible, so βNoβ is an acceptable answer.
-Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
-β’
-The instructions should contain the exact command and environment needed to run to reproduce the results.
-See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
-β’
-The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
-β’
-The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines.
-If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
-β’
-At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
-β’
-Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
-β’
-6.
-Experimental setting/details
-Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
-Answer: [Yes]
-Guidelines:
-β’
-The answer NA means that the paper does not include experiments.
-β’
-The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
-β’
-The full details can be provided either with the code, in appendix, or as supplemental material.
-β’
-7.
-Experiment statistical significance
-Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
-Answer: [No]
-Justification: We do not report error bars following prior arts [jing2021hinet, zhu2018hidden, ke2024stegformer, baluja2017hiding, lu2021large].
-Guidelines:
-β’
-The answer NA means that the paper does not include experiments.
-β’
-The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
-β’
-The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
-β’
-The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
-β’
-The assumptions made should be given (e.g., Normally distributed errors).
-β’
-It should be clear whether the error bar is the standard deviation or the standard error of the mean.
-β’
-It is OK to report 1-sigma error bars, but one should state it.
-The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
-β’
-For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
-β’
-If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
-β’
-8.
-Experiments compute resources
-Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
-Answer: [Yes]
-Justification: The training costs are reported in Appendix Section B.1.
-Guidelines:
-β’
-The answer NA means that the paper does not include experiments.
-β’
-The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
-β’
-The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
-β’
-The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didnβt make it into the paper).
-β’
-9.
-Code of ethics
-Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
-Answer: [Yes]
-Justification: The research in the paper conform with the NeurIPS Code of Ethics.
-Guidelines:
-β’
-The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
-β’
-If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
-β’
-The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
-β’
-10.
-Broader impacts
-Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
-Answer: [Yes]
-Justification: The broader impacts have been discussed in Appendix Section H.
-Guidelines:
-β’
-The answer NA means that there is no societal impact of the work performed.
-β’
-If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
-β’
-Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
-β’
-The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments.
-However, if there is a direct path to any negative applications, the authors should point it out.
-For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation.
-On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
-β’
-The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
-β’
-If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
-β’
-11.
-Safeguards
-Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
-Answer: [N/A]
-Justification: The paper poses no such risks.
-Guidelines:
-β’
-The answer NA means that the paper poses no such risks.
-β’
-Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
-β’
-Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
-β’
-We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
-β’
-12.
-Licenses for existing assets
-Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
-Answer: [Yes]
-Justification: Please refer to Appendix Section A.
-Guidelines:
-β’
-The answer NA means that the paper does not use existing assets.
-β’
-The authors should cite the original paper that produced the code package or dataset.
-β’
-The authors should state which version of the asset is used and, if possible, include a URL.
-β’
-The name of the license (e.g., CC-BY 4.0) should be included for each asset.
-β’
-For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
-β’
-If assets are released, the license, copyright information, and terms of use in the package should be provided.
-For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets.
-Their licensing guide can help determine the license of a dataset.
-β’
-For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
-β’
-If this information is not available online, the authors are encouraged to reach out to the assetβs creators.
-β’
-13.
-New assets
-Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
-Answer: [N/A]
-Justification: This paper does not release new assets.
-Guidelines:
-β’
-The answer NA means that the paper does not release new assets.
-β’
-Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
-β’
-The paper should discuss whether and how consent was obtained from people whose asset is used.
-β’
-At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
-β’
-14.
-Crowdsourcing and research with human subjects
-Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
-Answer: [N/A]
-Justification: This paper does not involve crowdsourcing nor research with human subjects.
-Guidelines:
-β’
-The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
-β’
-Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
-β’
-According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
-β’
-15.
-Institutional review board (IRB) approvals or equivalent for research with human subjects
-Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
-Answer: [N/A]
-Justification: This paper does not involve crowdsourcing nor research with human subjects.
-Guidelines:
-β’
-The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
-β’
-Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research.
-If you obtained IRB approval, you should clearly state this in the paper.
-β’
-We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
-β’
-For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
-β’
-16.
-Declaration of LLM usage
-Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research?
-Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
-Answer: [Yes]
-Guidelines:
-β’
-The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
-β’
-Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.
-β’
+13.4 Potential Risks and Mitigation Strategies
+At the same time, it is important to recognize the dual-use nature of steganography: while it can be applied to strengthen privacy and protect intellectual property, it may also pose risks if used for malicious concealment of content.
+Specifically, the ability to embed large-scale semantic content in natural images could be exploited for covert communication or data exfiltration, bypassing conventional content moderation or security inspection systems.
+To mitigate such risks, we emphasize that our research is conducted solely for academic purposes, aiming to enhance the understanding of multi-modal information hiding and detection.
+We encourage future work to develop complementary detection and forensic tools that can identify and regulate improper applications of this technology.
+14 Background and Preliminaries
+14.1 Terminology
+Steganography.
+Steganography is a technique for secret communication, which aims to hiding secret information in other, unsuspected data.
+As defined in [24, 2], Steganographic methods generally do rely on the assumption that the existence of the covert communication is unknown to third parties and are mainly used in secret point-to-point communication between trusting parties.
+As a result, steganographic methods are in general not robust, i.e., the hidden information cannot be recovered after data manipulation.
+In a word, the goal of steganography is to communicate securely in a completely undetectable manner and to avoid drawing suspicion to the transmission of a hidden message, not the robust.
+Watermark.
+As opposed to steganography, watermark has the additional notion of robustness against attacks [24].
+A practical implication of the robustness requirement is that watermarking methods can typically embed much less information into host data than steganographic methods.
+However, steganography and watermarking are thus more complementary than competitive approaches.
+In the experiment, we remove the all robustness requirement of StegaStamp [60] and treat it as a steganographic model.
+However, for an overview of steganography and watermark the reader is referred to [24], [59], and [51].
+Cover and Stego.
+In steganography, the cover refers to the original, unmodified object used as the carrier for hidden information [19].
+After embedding the secret message, the resulting object is called the stego.
+Ideally, the stego object should remain indistinguishable from the cover object, ensuring that the hidden information cannot be detected by human observers or automated steganalysis models.
+15 Additional Results